### v2

# A Second Main Theorem for Entire Curves Intersecting Three Conics

Abstract.
We establish a Second Main Theorem for entire holomorphic curves $f:\mathbb{C}\to\mathbb{P}^{2}$ intersecting a generic configuration of three conics $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$ in the complex projective plane $\mathbb{P}^{2}$. Using invariant logarithmic $2$-jet differentials with negative twists, we prove the estimate
$T_{f}(r)\leqslant 5\sum_{i=1}^{3}N_{f}^{[1]}(r,\mathcal{C}_{i})+o\big(T_{f}(r)\big)\quad\parallel,$
where $T_{f}(r)$ is the Nevanlinna characteristic function, and $N_{f}^{[1]}(r,\mathcal{C}_{i})$ is the $1$-truncated counting function.
The key innovation of our approach is establishing new vanishing lemmas of the form
$H^{0}\bigl(\mathbb{P}^{2},\,E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\bigr)=0$
for specific pairs $(m,t)$, achieved by combining algebro-geometric arguments with computer-assisted computations through a mod-$p$ reduction technique. This yields a systematic method for proving vanishing results for negatively twisted jet differentials—a key component in complex hyperbolic geometry.
Key words and phrases:
Nevanlinna theory, Second Main Theorem, entire curves, invariant logarithmic $2$-jet differentials, Siu’s slanted vector fields, Demailly-Semple tower
2020 Mathematics Subject Classification:
32H30, 32Q45
Appendix B was authored by Pengchao Wang.
1. Introduction
1.1. Motivation
In the complex projective plane $\mathbb{P}^{2}$, the logarithmic Kobayashi conjecture asserts that if $\mathcal{C}\subset\mathbb{P}^{2}$ is a generic algebraic curve of degree $d\geqslant 5$, then the complement $\mathbb{P}^{2}\setminus\mathcal{C}$ should be hyperbolic [Kobayashi1970]. Various notions of hyperbolicity exist; here we use the most straightforward one: $\mathbb{P}^{2}\setminus\mathcal{C}$ is said to be hyperbolic if there is no nonconstant holomorphic map $f:\mathbb{C}\to\mathbb{P}^{2}\setminus\mathcal{C}$. This conjecture generalizes the classical Little Picard Theorem on the hyperbolicity of $\mathbb{P}^{1}\setminus\{\text{3 points}\}$, and is supported by explicit examples constructed by Zaidenberg [Zaidenberg1988]. For generic curves $\mathcal{C}$ of sufficiently large degree in $\mathbb{P}^{2}$, the conjecture was confirmed by Siu and Yeung [siu_yeung1996] by constructing explicit logarithmic $2$-jet differentials with negative twists. Later, building on the work [Demailly-Elgoul2000], El Goul [Elgoul2003] improved the degree bound to $15$ for very generic $\mathcal{C}$. Combining this with Siu’s strategy of using slanted vector fields [Siu2004] and a technique of Păun [mihaipaun2008], Rousseau [Rousseau2009] further reduced the degree bound to $14$.
These results for high-degree curves naturally extend to configurations with multiple components. In the case where the curve $\mathcal{C}=\cup_{i=1}^{q}\mathcal{C}_{i}$ decomposes into the union of $q\geqslant 2$ generic algebraic curves $\mathcal{C}_{i}$ of degree $d_{i}$, the hyperbolicity of the complement $\mathbb{P}^{2}\setminus\mathcal{C}$ has been established with smaller degree bounds $d=\sum_{i=1}^{q}d_{i}$ in various works. For instance, a result in [babets1984] settles the case where $q\geqslant 5$. The case $(q=4,\sum_{i=1}^{4}d_{i}\geqslant 5)$ was treated in [DSW1995-AJM]. The configuration of three conics was considered in [Grauert1989, DSW1995-AJM, Francois-Duval-2001] by different methods. Dethloff, Schumacher, and Wong [DSW1995-duke] employed the symmetric differentials $1$-forms to deal with the case of two conics and one cubic. Rousseau [Rousseau2009] addressed the case $q=2$ under the assumptions $(d_{1}=1,d_{2}\geqslant 11)$, $(d_{1}=2,d_{2}\geqslant 8)$, $(d_{1}=3,d_{2}\geqslant 5)$, or $(d_{1}=4,d_{2}\geqslant 4)$. Notably, for $q\geqslant 3$, Noguchi, Winkelmann, and Yamanoi [NWA2007] established the algebraic degeneracy of every entire curve in $\mathbb{P}^{2}\setminus\mathcal{C}$ under the mild condition $\sum_{i=1}^{q}d_{i}\geqslant 4$. Their result also extends to the corresponding higher-dimensional setting.
Thus, in the aforementioned settings, every algebraically nondegenerate entire curve $f:\mathbb{C}\rightarrow\mathbb{P}^{2}$ must intersect the curve $\mathcal{C}$. In this article, we aim to quantify these geometric results by means of Nevanlinna theory. Specifically, we aim to establish a Second Main Theorem type estimate:
(1.1) $T_{f}(r)\leqslant c\sum_{i=1}^{q}N_{f}^{[k]}(r,\mathcal{C}_{i})+o\big(T_{f}(r)\big)\quad\parallel,$
where
$N^{[k]}(r,\mathcal{C}_{i}):=\int_{1}^{r}\sum_{|z|<t}\min\{k,\operatorname{ord}_{z}f^{*}\mathcal{C}_{i}\}\,\frac{\operatorname{d}t}{t}\quad(i=1,\dots,q;\,r>1)$
are the $k$-truncated counting functions, measuring the frequency of intersections $f(\mathbb{C})\cap\mathcal{C}_{i}$; and
$T_{f}(r):=\int_{1}^{r}\frac{\operatorname{d}t}{t}\int_{|z|<t}f^{*}\omega_{\text{FS}}\quad(r>1)$
is the order function, capturing the area growth rate of the holomorphic discs $f(\{|z|<t\})$ with respect to the Fubini-Study metric $\omega_{\text{FS}}$ on $\mathbb{P}^{2}$. The constant $c$ on the right-hand side of (1.1) is independent of $r$, and the notation $\parallel$” indicates that the inequality (1.1) holds for all $r>1$ outside a set of finite Lebesgue measure.
When $\mathcal{C}$ has at least $4$ irreducible components, such a Second Main Theorem type estimate with truncation at level $2$ was established by H. Cartan [Cartan1933] in the linear case where $\mathcal{C}_{i}$ are lines:
$T_{f}(r)\leqslant\frac{1}{q-3}\sum_{i=1}^{q}N_{f}^{[2]}(r,\mathcal{C}_{i})+o\big(T_{f}(r)\big)\quad\parallel.$
In the nonlinear case, analogous results (without truncation) were obtained by Eremenko and Sodin [Eremenko-Sodin1992] via pluripotential theory, and by Ru [Minru2009] through a synthesis of Cartan’s method and Diophantine geometry. These results also extend to the corresponding higher-dimensional setting. For further background and recent advances in Nevanlinna theory, see [Noguchi-Winkelmann2014, Ru21].
In the prior work [CHSX2025], the authors established a Second Main Theorem for $n+1$ hypersurfaces $\{D_{i}\}_{i=1}^{n+1}$ in $\mathbb{P}^{n}$ in general position and intersecting transversally, proving the defect relation
(1.2) $\sum_{i=1}^{n+1}\delta_{f}(D_{i})<n+1$
for any algebraically nondegenerate entire curve $f:\mathbb{\mathbb{C}}\rightarrow\mathbb{P}^{n}$ under the optimal total degree assumption $\sum_{i=1}^{n+1}\deg D_{i}\geqslant n+2$. Here, each defect $\delta_{f}(D_{i})$ is defined by
$\delta_{f}(D_{i})\coloneqq 1-\limsup_{r\to\infty}\frac{N_{f}(r,D_{i})}{\deg(D_{i})\,T_{f}(r)}.$
The proof combined parabolic Nevanlinna theory [Siboni-Paun2014] with the Second Main Theorems for semi-Abelian varieties [Noguchi-Yamanoi-Winkelmann2002] and for $n+2$ hypersurfaces [Ru2004]. Subsequently, Ru and Wang [Ru-Wang2025] extended the result (1.2) to moving targets with truncation level $1$, while preserving the upper bound $n+1$ on the right-hand side.
In the case of $n=2$, the defect relation (1.2) might be viewed as the weakest, albeit nontrivial, form of a Second Main Theorem in the vein of (1.1) with noneffective $c<\infty$. The primary objective of this article is to establish an effective Second Main Theorem, with truncation level $1$, for three generic conics with an explicit constant $c=5$.
1.2. Two Key Theorems
The pursuit of a Second Main Theorem encounters a fundamental obstacle: the very nature of entire curves exhibits an inherent and wild flexibility in large-scale geometry, as evidenced by the exotic examples in [HX2021, Xie2024, WX2025, CFX2025]. Our strategy constrains this flexibility by leveraging additional algebro-geometric structures to extract the underlying rigidity of entire curves, building on the following two key results.
Theorem 1.1 (McQuillan [Mcquillan1998]).
Let $\mathcal{C}=\sum_{i=1}^{q}\mathcal{C}_{i}$ be a normal crossings divisor in $\mathbb{P}^{2}$, where each $\mathcal{C}_{i}$ is a smooth algebraic curve of degree $d_{i}$ with $d=\sum_{i=1}^{q}d_{i}\geqslant 4$. Let $f:\mathbb{C}\rightarrow\mathbb{P}^{2}$ be an algebraically nondegenerate entire holomorphic curve. If the lifting $f_{[1]}:\mathbb{C}\rightarrow X_{1}=\mathbb{P}\,T_{\mathbb{P}^{2}}(-\mathrm{log}\,\mathcal{C})$ is algebraically degenerate, then the following Second Main Theorem type estimate holds:
$(d-3)\,T_{f}(r)\leqslant N^{[1]}_{f}(r,\mathcal{C})+o\big(T_{f}(r)\big)\quad\parallel.$
Theorem 1.2 ([Huynh-Vu-Xie-2017, Theorem 3.1]).
Let $\mathcal{C}=\sum_{i=1}^{q}\mathcal{C}_{i}$ be a normal crossings divisor in $\mathbb{P}^{2}$. For any logarithmic jet differential
$\omega\in H^{0}\big(\mathbb{P}^{2},E_{k,m}^{GG}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)\quad(k,m,t\geqslant 1),$
and any entire holomorphic curve $f:\mathbb{C}\rightarrow\mathbb{P}^{2}$ such that $f^{*}\omega\not\equiv 0$, the following Second Main Theorem type estimate holds:
(1.3) $T_{f}(r)\leqslant\frac{m}{t}\sum_{i=1}^{q}N_{f}^{[1]}(r,\mathcal{C}_{i})+O\big(\mathrm{log}\,T_{f}(r)+\mathrm{log}\,r\big)\quad\parallel.$
Here, $E_{k,m}^{GG}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$ denotes the logarithmic Green–Griffiths $k$-jet bundle of weighted order $m$ (see Section 2). The use of such higher-order jet differentials in the study of complex hyperbolicity goes back to André Bloch [Bloch1926]. In the present work, we restrict attention to the “invariant” subbundle $E_{k,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\subset E_{k,m}^{GG}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$, originally introduced by Demailly [Demailly1997]. This subbundle possesses stronger positivity properties, making it particularly suitable for constructing sections; further details can be found in Section 2.
Theorem 1.2 can be seen as a quantitative refinement of the Fundamental Vanishing Theorem for entire curves (cf. [Green-Griffiths1980, pp. 55-59], [Siu-Yeung1996, Lemma 1.2] and [Demailly1997, Corollary 7.9]). Although the original statement in [Huynh-Vu-Xie-2017] is formulated for hypersurfaces in $\mathbb{P}^{n}$, the proof extends directly to the case of simple normal crossing divisors in projective manifolds (cf. [Ru21, Theorem A10.7.2]).
1.3. Our First Result
The first result of this article is an application of Theorem 1.2 in the case $k=1$.
Theorem 1.3.
Let $\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i}$ be a union of three smooth curves $\mathcal{C}_{i}$ of degrees $d_{i}$ with simple normal crossings. Assuming that $d_{1}\geqslant d_{2}\geqslant d_{3}\geqslant 2$ and $d_{1}\geqslant 3$, so that the total degree $d=d_{1}+d_{2}+d_{3}\geqslant 7$. For a nonconstant entire holomorphic curve $f\colon\mathbb{C}\rightarrow\mathbb{P}^{2}$ whose image is not contained in any $\mathcal{C}_{i}$, one has the estimate
$T_{f}(r)\leqslant c\sum_{i=1}^{3}N_{f}^{[1]}(r,\mathcal{C}_{i})+o\big(T_{f}(r)\big)\quad\parallel,$
for any positive constant $c$ satisfying
$c>\frac{1}{(d-3)\delta_{1}},\quad\text{where }\quad{\delta_{1}=\tfrac{1}{2}-\tfrac{\sqrt{9(d-1)^{2}-12(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})}}{6(d-3)}}.$
Here, the numerical assumption on $d_{1},d_{2},d_{3}$ is made to guarantee that the Euler characteristics of the logarithmic symmetric differential bundles $S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathscr{C})$ are positive:
$\chi\big(S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathscr{C})\big)=O(1)^{+}\cdot m^{3}+O(m^{2})>0,\quad\forall\,m\gg 1.$
Remark 1.4.
In the case where $\mathcal{C}$ consists of two conics and one cubic, a direct computation shows that $\delta_{1}$ takes the value $\frac{6-\sqrt{33}}{12}$. This leads to a Second Main Theorem of the form
$T_{f}(r)\leqslant c\sum_{i=1}^{3}N_{f}^{[1]}(r,\mathcal{C}_{i})+o\big(T_{f}(r)\big)\quad\parallel,$
valid for any positive constant $c$ satisfying $c>\frac{3}{6-\sqrt{33}}\approx 11.74\ldots$. This quantifies the hyperbolicity result of $\mathbb{P}^{2}\setminus\mathcal{C}$ due to Dethloff–Schumacher–Wong [DSW1995-duke].
1.4. Proof of Theorem 1.3
First, if the image of $f$ is contained in an irreducible curve $\mathcal{D}$, then $\mathcal{D}$ must intersect the configuration $\{\mathcal{C}_{i}\}$ in at least three points [DSW1995-AJM, Prop. 4.1]. In this special case, stronger results are already available in the literature [Ru21, Chapter 4].
We now assume that $f$ is algebraically nondegenerate. The key step is to find a nonzero global negatively twisted logarithmic symmetric differential form
(1.4) $\omega\in H^{0}\big(\mathbb{P}^{2},\,S^{m}\,T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}(-t)\big),\quad m\geqslant 1,\ t\geqslant 1,\ m/t\leqslant c,\ \omega\not\equiv 0.$
Proof of Theorem 1.3 assuming (1.4).
If $f^{*}\omega\not\equiv 0$, then Theorem 1.2 yields the desired estimate. On the other hand, if $f^{*}\omega\equiv 0$, then we can apply McQuillan’s Theorem 1.1 to obtain the stronger inequality
$(d-3)\,T_{f}(r)\leqslant N_{f}^{[1]}(r,\mathcal{C})+o(T_{f}(r))\quad\parallel.$
In either case, the conclusion of the theorem holds. ∎
The proof of (1.4) will be given in Section 3, based on Riemann–Roch computations and a Bogomolov-type vanishing theorem (Theorem 2.1).
1.5. Our Second, Main Result: Second Main Theorem for Three Conics.
Having established an effective Second Main Theorem for configurations of large degree, we now turn to the more challenging case of three generic conics $\mathcal{C}$ in $\mathbb{P}^{2}$. While the $1$-jet method proves successful in high-degree regimes, it encounters a fundamental obstruction in this low-degree setting, as the Euler characteristics are now negative:
$\chi\big(S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathscr{C})\big)=O(1)^{-}\cdot m^{3}+O(m^{2})<0,\quad\forall\,m\gg 1.$
Question 1.5.
For a generic configuration of three conics $\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i}$ in $\mathbb{P}^{2}$, does there exist some integer $m\geqslant 1$ such that
$H^{0}\bigl(\mathbb{P}^{2},\,S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-1)\bigr)\neq 0?$
Here and throughout, “generic” means outside a proper Zariski-closed subset of the parameter space.
It is instructive to compare this with known borderline cases. On the one hand, if we replace the three conics $\mathcal{C}$ by certain special six lines, such an $m$ does exist (see [Hirzebruch, p. 133] and also [VZ2000, Theorem 1.4]). On the other hand, if $\mathcal{C}$ is a generic sextic curve, then no such $m$ exists [Elgoul2003, Lemma 1.4.1].
Since the answer to Question 1.5 remains unknown, [^1] Using the approach developed in this work, we can verify the vanishing for $1\leqslant m\leqslant 7$ for generic three conics. the key assumption (1.4) cannot be verified a priori. This situation illustrates the principle that “the lower the degree, the harder the hyperbolicity” [Duval2021, p. 3]. To circumvent this obstacle, we employ invariant logarithmic $2$-jet differentials with negative twists. The main result of this article is as follows.
Theorem 1.6 (Second Main Theorem for Three Conics).
For a generic configuration of three conics $\{\mathcal{C}_{i}\}_{i=1}^{3}$ in $\mathbb{P}^{2}$ and a nonconstant entire holomorphic curve $f\colon\mathbb{C}\to\mathbb{P}^{2}$ whose image is not contained in any $\mathcal{C}_{i}$, the following inequality holds:
(1.5) $T_{f}(r)\leqslant 5\sum_{i=1}^{3}N_{f}^{[1]}(r,\mathcal{C}_{i})+o\big(T_{f}(r)\big)\quad\parallel.$
Remark 1.7.
The constant $c=5$ in this theorem is smaller than the constant ${\approx}11.74$ appearing in Theorem 1.3 for the case of two conics and one cubic. This may seem counterintuitive at first, since the smaller total degree $2+2+2<2+2+3$ would generally suggest a larger constant. The explanation lies in the method of proof: here we employ not only $1$-jet technique but also the more refined $2$-jet approach. A detailed analysis using the latter is expected to yield sharper constants not only for the two-conics-and-one-cubic case but also for other configurations covered by Theorem 1.3.
Remark 1.8.
An unexpected outcome of our method is that, as a byproduct, it allows us to explicitly identify certain configurations of three Fermat-type conics $\mathcal{C}$ satisfying the Second Main Theorem type estimate in (1.5) via an alternative “direct” argument. For instance, consider the three conics $\mathcal{C}$ defined by the quadratic polynomials:
$2Z_{0}^{2}+Z_{1}^{2}+Z_{2}^{2},\qquad Z_{0}^{2}+2Z_{1}^{2}+Z_{2}^{2},\qquad Z_{0}^{2}+Z_{1}^{2}+2Z_{2}^{2}.$
Further details can be found in Remark 7.5. This also explains why we are satisfied with the constant $5$ in the Main Theorem, rather than attempting to lower it slightly by proving additional vanishing results—which, while manageable, are not very interesting.
We now outline the main ideas underlying the proof of Theorem 1.6 for a finite constant $c<\infty$, before refining the argument to achieve the sharper goal $c=5$. As in the proof of Theorem 1.3, we can assume that $f$ is algebraically nondegenerate.
The core idea of the proof is to seek two “independent”, negatively twisted invariant logarithmic $2$-jet differentials $\omega_{1}$ and $\omega_{2}$ such that:
  • If $f^{*}\omega_{1}\not\equiv 0$ or $f^{*}\omega_{2}\not\equiv 0$, then Theorem 1.2 directly implies a desired Second Main Theorem type estimate.
  • If $f^{*}\omega_{1}\equiv 0$ and $f^{*}\omega_{2}\equiv 0$, then the “independence” of $\omega_{1}$ and $\omega_{2}$ ensures that the lifting $f_{[1]}(\mathbb{C})\subset X_{1}$ is algebraically degenerate, where $X_{1}$ is the first level of the logarithmic Demailly-Semple tower (see Subsection 2.2) for $(\mathbb{P}^{2},\mathcal{C})$. Hence McQuillan’s Theorem 1.1 applies.
In either case, we obtain the desired Second Main Theorem type estimate.
In practice, however, this conceptual framework faces substantial challenges. Controlling the common base locus of $\omega_{1}$ and $\omega_{2}$—essential for ensuring their independence—requires detailed geometric information that is typically accessible only through explicit constructions of negatively twisted logarithmic jet differentials. While such constructions are feasible when the curve $\mathcal{C}$ is of high degree, as demonstrated by Siu and Yeung [siu_yeung1996], the low-degree case presents significantly greater difficulties. To date, not a single such differential has been explicitly described for low-degree curves, due to the intricate algebraic complexity of jet differentials (cf. [Merker2008]). Moreover, the subsequent elimination step introduces further difficulties, rendering a direct approach effectively intractable.
A notable illustration of these difficulties arises even in the $1$-jet case. Debarre [Debarre2005] conjectured that for a generic choice of high-degree hypersurfaces $H_{1},\dots,H_{c}\subset\mathbb{P}^{N}$ with $c\geqslant N/2$, the complete intersection $X:=H_{1}\cap\cdots\cap H_{c}$ should have an ample cotangent bundle $T_{X}^{*}$, which is equivalent to the existence of sufficiently many negatively twisted $1$-jet differentials on $X$ with empty common base locus. The first proof of Debarre’s ampleness conjecture was established by Xie [Xie2018], using explicit $1$-jet differentials obtained by Brotbek [Brotbek2016Symmetricdifferentialforms] for Fermat-type hypersurfaces. A key step in Xie’s approach is a core lemma [Xie2018, p. 969] that overcomes substantial difficulties in the elimination process. Shortly thereafter, Brotbek and Darondeau [BD2018] found an alternative proof, which instead applies Nakamaye’s theorem [Nakamaye] to tackle the elimination.
1.6. Roadmap for Theorem 1.6: Towards an Effective Finite Constant $c<\infty$
We now outline the proof of our main result, focusing first on establishing an explicit finite constant $c$ in the inequality (1.5). The argument proceeds in several steps.
First, for generic three conics $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$, standard Riemann–Roch computations (detailed in Section 4) guarantee the existence of an irreducible, negatively twisted invariant logarithmic $2$-jet differential
(1.6) $\omega_{1}\in H^{0}\big(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big),$
with weighted degree $m\geqslant 1$, vanishing order $t\geqslant 1$.
If $f^{*}\omega_{1}\not\equiv 0$, then Theorem 1.2 immediately yields the desired Second Main Theorem with a finite constant $c$. We may therefore assume $f^{*}\omega_{1}\equiv 0$ in what follows.
The subsequent strategy depends on the values of $(m,t)$:
Case 1: $t>7$. We apply Siu’s strategy [Siu2004] of slanted vector fields, which realized in dimension two by Paŭn [mihaipaun2008] (compact case) and Rousseau [Rousseau2009] (log case), to differentiate $\omega_{1}$ by a slanted vector field with twist $\mathcal{O}_{\mathbb{P}^{2}}(7)$, thereby obtaining an “independent” nonzero invariant logarithmic $2$-jet differential
$\omega_{2}\in H^{0}\big(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t+7)\big).$
Consequently, the proof can be completed by standard arguments (see Section 5).
Case 2:  $t<m$. Following an approach of Demailly–El Goul [Demailly-Elgoul2000] (see Section 6), we can show the existence of a nonzero invariant logarithmic $2$-jet differential $\omega_{2}$ with negative twist via Riemann-Roch computations. However, the definition of $\omega_{2}$ is subtle: it is not defined on the whole space $\mathbb{P}^{2}$, but only on the zero locus $\{\omega_{1}=0\}$ within the second level $X_{2}$ of the Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C})$; see Proposition 6.2. We then distinguish two subcases:
  • If $f^{*}\omega_{2}\not\equiv 0$, we can derive a Second Main Theorem-type estimate; see Theorem 6.4. The arguments are analogous to those for Theorem 1.2 but are refined in a crucial way, using the structure of Wronskians (see Proposition 6.3). This refinement represents a new contribution of this article.
  • If $f^{*}\omega_{2}\equiv 0$, then the simultaneous vanishing of $f^{*}\omega_{1}$ and $f^{*}\omega_{2}$ implies that the lifting $f_{[1]}(\mathbb{C})$ is algebraically degenerate in the first level $X_{1}$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C})$, and hence McQuillan’s Theorem 1.1 applies.
Case 3: The Remaining Possibilities. For the remaining pairs $(m,t)$ not covered by the previous two cases, we establish the following Key Vanishing Lemma I to rule out the existence of any nonzero $\omega_{1}$ in these settings. Note that vanishing for $(m,t)$ implies vanishing for all $(m,t^{\prime})$ with $t^{\prime}>t$. Thus we complete the proof of a Second Main Theorem for three generic conics with a finite constant $c$. Moreover, careful numerical computation shows that we can take $c=19$; see Subsection 6.4.
1.7. New Vanishing Results and Proof Strategy
Key Vanishing Lemma I.
For a generic configuration of three conics $\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i}$ in $\mathbb{P}^{2}$, one has:
$H^{0}\big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)=0$
for all $(m,t)$ in the set $\{(3,3),(4,4),(5,5),(6,6),(7,7)\}.$
While such vanishing results have long been recognized as fundamental in complex hyperbolicity theory, a proof has remained out of reach for lack of a viable method. The principal innovation of this article is a novel methodological framework that resolves this impasse.
Our proof of Key Vanishing Lemma I proceeds in three steps, presented in Subsection 7.2:
  • Algebraic Reduction: We begin by transforming the problem of constructing invariant logarithmic jet differentials into that of solving a specific, tractable linear system, leveraging the gaga principle to apply algebraic techniques in an analytic setting. The central challenge then becomes establishing bounds on the exponents in the denominators of the associated coefficients. This challenge is addressed by a careful choice of basis and an analysis of the transition matrices between different frames.
  • Computational Verification: For a set of well-chosen explicit conics, we provide a computer-assisted proof that the associated linear systems possess only the trivial zero solution, as verified via symbolic computation.
  • Geometric Extension: A semi-continuity argument globalizes the computationally verified vanishing results from a specific choice of three conics to generic ones, concluding the proof.
Remark 1.9.
An alternative approach to establishing such vanishing results is through cohomological diagram chasing based on the filtration structure (2.3). This method enables systematic construction of long exact sequences in cohomology, thereby facilitating stepwise computation of $H^{0}\big(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}(-t)\big)$. When $m=t$ is not divisible by $3$, the vanishing can be reduced to that of certain logarithmic $1$-jet differentials with negative twists. This observation is due to Pengchao Wang; see Appendix B.222 If the answer to Question 1.5 is affirmative, McQuillan’s Theorem 1.1 immediately yields an effective second main theorem for three generic conics. Should the answer be negative, Proposition B.2 automatically establishes the cases $(m,t)=(4,4),(5,5),(7,7)$ in Key Vanishing Lemma I.
However, in more general cases—such as when $m>t$ or when $3$ divides m=tm=t—this diagram-chasing approach becomes substantially more complicated and its feasibility remains uncertain. Determining the essential information required for the diagram chasing presents a significant challenge. In particular, computing $H^{i}$ for $i=0,1,2$ for each term in the long exact sequence requires either analyzing the connecting homomorphisms from $H^{i}$ to $H^{i+1}$, or computing other $H^{i}$ groups that typically do not vanish, thereby introducing considerable additional complexity.
Remark 1.10.
To demonstrate the efficacy of our approach, we apply the systematic framework developed in this article to establish new vanishing results about $2$-jet differentials, which in turn yield improved degree bounds for the Kobayashi hyperbolicity conjecture in complex dimension two. Specifically, we refine prior results by lowering Păun’s [mihaipaun2008] bound from $18$ to $16$ for very generic surfaces in $\mathbb{P}^{3}$, and Rousseau’s [Rousseau2009] bound from $14$ to $12$ for complements of generic curves in $\mathbb{P}^{2}$. Complete details will be presented in a forthcoming paper.
1.8. Proof Blueprint of Theorem 1.6 with $c=5$
We first show that in the existence result for $\omega_{1}$ in (1.6), the ratio $t/m$ can be made close to $\frac{4-\sqrt{10}}{3}$; see Proposition 4.2. If $f^{*}\omega_{1}\not\equiv 0$, then Theorem 1.2 yields the estimate (1.5) with constant $c$ around $\frac{3}{4-\sqrt{10}}=3.58\ldots$. This value represents the theoretical limit of our current approach. Our goal of achieving $c=5$ is strategically set near this limit.
In Case 1 ($t>7$), to ensure that the application of Theorem 1.2 to both $\omega_{1}$ and $\omega_{2}$ yields $c=5$, we strengthen the conditions to $m/t\leqslant 5$ and $m/(t-7)\leqslant 5$, which is equivalent to $t\geqslant m/5+7$.
In Case 2 ($t<m$), we refine Demailly–El Goul’s approach [Demailly-Elgoul2000] to estimate the ratio of the vanishing order of $\omega_{2}$ to its weighted degree. Through a detailed and refined analysis in Section 6, we show that the constant $c$ can be expressed as $2/\tau_{1}(\frac{t}{m})$, where
$\tau_{1}(\tfrac{t}{m})=\frac{3}{2}\left(4-\frac{t}{m}\right)-\frac{\sqrt{3}}{2}\sqrt{32-8\frac{t}{m}+3\left(\frac{t}{m}\right)^{2}}.$
A direct computation shows that $2/\tau_{1}(\frac{t}{m})\leqslant 5$ is equivalent to $t\leqslant\frac{92}{135}m$.
Although most positive integer pairs $(m,t)$ satisfy either $t\leqslant\frac{92}{135}m$ or $t\geqslant m/5+7$, exceptional cases remain where neither condition holds. To address these, we need the following new vanishing results:
Key Vanishing Lemma II.
For a generic configuration of three conics $\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i}$ in $\mathbb{P}^{2}$, one has:
(1.7) $H^{0}\big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)=0$
for all $(m,t)\in\{(3,3),(4,3),(5,4),(6,5),(7,5),(8,6),(9,7),(10,7),(11,8),(12,9),(13,9)\}$.
The proof (see Subsection 7.3) introduces two new ingredients: the utilization of Fermat-type conics to facilitate computation; and a modulo $p=5$ technique that leverages computer effectively perform Gaussian elimination, thus proving that the associated system of linear equations with integer coefficients has only the trivial solution.
1.9. Structure of the article
Section 2 provides necessary background. Section 3 presents detailed Riemann–Roch computations that form the core of the proof of Theorem 1.3. Section 4 advances to more sophisticated Riemann–Roch computations, specifically designed to establish the existence of negatively twisted invariant logarithmic $2$-jet differentials for three generic conics. Section 5 details the application of Siu’s slanted vector fields in Case 1. Section 6 then introduces a refinement of the Demailly–El Goul argument, adapted to handle Case 2. The main innovation of the article is contained in Section 7, where we overcome fundamental obstructions in our low-degree setting by establishing Key Vanishing Lemmas. Finally, Appendix A provides the Maple code used to verify these vanishing results, and Appendix B presents an observation by Pengchao Wang.
2. Preliminaries
2.1. Logarithmic Jet Bundles and Invariant Differential Forms
Let $X$ be an $n$-dimensional complex manifold. The $k$-th jet bundle $J_{k}(X)$ is defined by considering holomorphic germs $f\colon(\mathbb{C},0)\to(X,x)$, where two germs are considered equivalent if their Taylor expansions coincide up to order $k$ in some local coordinate system. The equivalence class of a germ $f$, denoted by $j_{k}(f)$, is said to be regular if $f^{\prime}(0)\neq 0$.
The total space $J_{k}(X)=\cup_{x\in X}J_{k}(X)_{x}$ carries a natural topology and, together with the natural projection $\pi_{k}\colon J_{k}(X)\to X$, forms a holomorphic fiber bundle. Note that $J_{1}(X)$ coincides with the tangent bundle $T_{X}$.
Now let $D\subset X$ be a normal crossing divisor. The logarithmic cotangent bundle $T_{X}^{*}(\mathrm{log}\,D)$ is a locally free sheaf generated by meromorphic $1$-forms with at most logarithmic poles along $D$. Precisely, for $x\in X\setminus D$, the stalk $T_{X,x}^{*}(\mathrm{log}\,D)$ coincides with the stalk $T_{X,x}^{*}$ of the usual cotangent sheaf $T_{X}^{*}$ at $x$. For $x\in D$, take a neighborhood $U$ of $x$ and irreducible holomorphic functions $\eta_{1},\dots,\eta_{\ell}$ on $U$ such that $D\cap U=\{\eta_{1}\cdots\eta_{\ell}=0\}$; then the stalk is given by $T_{X,x}^{*}(\mathrm{log}\,D)=\sum_{j=1}^{\ell}\mathcal{O}_{X,x}\frac{\mathrm{d}\eta_{j}}{\eta_{j}}+T_{X,x}^{*}$. Its dual, the logarithmic tangent bundle $T_{X}(-\mathrm{log}\,D)$, consists of holomorphic vector fields tangent to each component of $D$ (cf. [Noguchi1986, Elgoul2003]).
Consider an open subset $U\subset X$, a section $\omega\in H^{0}(U,T_{X}^{*}(\mathrm{log}\,D))$, and a jet $j_{k}(f)$ in $J_{k}(X)|_{U}$. The pullback $f^{*}\omega$ takes the form $A(z)\mathrm{d}z$ for some meromorphic function $A$, and the derivatives $A^{(j)}(0)$ for $0\leqslant j\leqslant k-1$ depend only on the jet $j_{k}(f)$. This construction induces a well-defined map:
(2.1) $\tilde{\omega}\colon J_{k}(X)|_{U}\to(\mathbb{C}\cup\{\infty\})^{k},\quad j_{k}(f)\mapsto\big(A(0),A^{\prime}(0),\dots,A^{(k-1)}(0)\big).$
A holomorphic section $s\in H^{0}(U,J_{k}(X))$ is called a logarithmic $k$-jet field on $U$ if for every open subset $U^{\prime}\subset U$ and every $\omega\in H^{0}(U^{\prime},T_{X}^{*}(\mathrm{log}\,D))$, the composition $\tilde{\omega}\circ s$ is holomorphic on $U^{\prime}$, where $\tilde{\omega}$ is defined as in (2.1). These logarithmic $k$-jet fields form a subsheaf of $J_{k}(X)$, which corresponds to the sections of a holomorphic fiber bundle over $X$ known as the logarithmic $k$-jet bundle along $D$, denoted $J_{k}(X,-\mathrm{log}\,D)$ (cf. [Noguchi1986, Elgoul2003, Rousseau2009]).
Furthermore, given a local basis $\omega_{1},\dots,\omega_{n}$ of $T_{X}^{*}(\mathrm{log}\,D)$ on $U$, we obtain a trivialization:
$H^{0}(U,J_{k}(X,-\mathrm{log}\,D))\to U\times(\mathbb{C}^{k})^{n},\quad\sigma\mapsto\big(\pi_{k}\circ\sigma;\;\tilde{\omega}_{1}\circ\sigma,\dots,\tilde{\omega}_{n}\circ\sigma\big).$
A logarithmic jet differential of order $k$ and weighted degree $m$ at a point $x\in X$ is a polynomial $Q(f^{(1)},\dots,f^{(k)})$ on the fiber of $J_{k}(X,-\mathrm{log}\,D)$ over $x$, satisfying the weighted homogeneity condition:
$Q\big(\lambda f^{\prime},\lambda^{2}f^{\prime\prime},\dots,\lambda^{k}f^{(k)}\big)=\lambda^{m}Q\big(f^{\prime},f^{\prime\prime},\dots,f^{(k)}\big),$
for all $\lambda\in\mathbb{C}^{*}$ and all $(f^{\prime},f^{\prime\prime},\dots,f^{(k)})\in J_{k}(X,-\mathrm{log}\,D)_{x}$.
The vector space $E_{k,m}^{GG}T_{X}^{*}(\mathrm{log}\,D)_{x}$ consists of all such jet differential polynomials defined at the point $x\in X$. Then the total space
$E_{k,m}^{GG}T_{X}^{*}(\mathrm{log}\,D):=\cup_{x\in X}E_{k,m}^{GG}T_{X}^{*}(\mathrm{log}\,D)_{x},$
acquires the structure of a holomorphic vector bundle over $X$ by Faà di Bruno’s formula (cf. e.g. [Constantine1996, Merker2015]), referred to as the logarithmic Green–Griffiths bundle of jet differentials of order $k$ and weighted degree $m$.
Let $\mathbb{G}_{k}$ denote the group of germs of $k$-jets of biholomorphisms of $(\mathbb{C},0)$:
$\mathbb{G}_{k}\coloneqq\Big\{\varphi\colon t\mapsto\sum_{j=1}^{k}c_{j}t^{j}\;\Big|\;c_{1}\in\mathbb{C}^{*},\;c_{j}\in\mathbb{C}\text{ for }j\geqslant 2\Big\},$
[…]
A logarithmic jet differential operator $P\in\Gamma\big(U,E_{k,m}^{GG}T_{X}^{*}(\mathrm{log}\,D)\big)$ over an open set $U\subset X$ is said to be invariant under the reparametrization group $\mathbb{G}_{k}$ if for every $\varphi\in\mathbb{G}_{k}$ the following identity holds:
(2.2) $P\big((f\circ\varphi)^{\prime},(f\circ\varphi)^{\prime\prime},\dots,(f\circ\varphi)^{(k)}\big)=(\varphi^{\prime})^{m}\cdot P\big(f^{\prime},f^{\prime\prime},\dots,f^{(k)}\big).$
The subbundle $E_{k,m}T_{X}^{*}(\mathrm{log}\,D)\subset E_{k,m}^{GG}T_{X}^{*}(\mathrm{log}\,D)$ is then defined as the holomorphic vector bundle whose fibers consist of all such invariant logarithmic jet differential polynomials.
This notion was initially introduced by Demailly [Demailly1997] in the compact case, and subsequently generalized to the logarithmic case by Dethloff-Lu [Dethloff-Lu2001].
Throughout this article, we only concern the situation $k=1$ or $2$. In the case that $\dim_{\mathbb{C}}X=2$, the bundle of invariant logarithmic jet differentials of order $2$ admits a filtration structure with the grading [Demailly1997]:
(2.3) $\mathrm{Gr}^{\bullet}E_{2,m}T_{X}^{*}(\mathrm{log}\,D)=\bigoplus_{0\leqslant j\leqslant\lfloor m/3\rfloor}S^{m-3j}T_{X}^{*}(\mathrm{log}\,D)\otimes\overline{K}_{X}^{j},$
where $\overline{K}_{X}$ is the logarithmic canonical line bundle $K_{X}\otimes\mathcal{O}_{X}(D)$. This filtration corresponds to the expression of an invariant logarithmic $2$-jet differential $P(f^{\prime},f^{\prime\prime})$ outside $D$ as:
$P(f^{\prime},f^{\prime\prime})=\sum_{0\leqslant j\leqslant\lfloor m/3\rfloor}\;\sum_{\begin{subarray}{c}\alpha_{1}+\alpha_{2}=m-3j\\ \alpha_{1},\alpha_{2}\geqslant 0\end{subarray}}R_{\alpha_{1},\alpha_{2},j}(f_{1},f_{2})\cdot(f_{1}^{\prime})^{\alpha_{1}}(f_{2}^{\prime})^{\alpha_{2}}(f_{1}^{\prime}f_{2}^{\prime\prime}-f_{2}^{\prime}f_{1}^{\prime\prime})^{j},$
where $f=(f_{1},f_{2})$ represents the holomorphic germs in local coordinates, and where $R_{\bullet}$ are holomorphic functions.
2.2. Logarithmic Demailly-Semple Tower
The main reference of this subsection is [Demailly1997, Dethloff-Lu2001].
For simplicity, consider the log-direct manifold $(X_{0},D_{0},V_{0})$ associated with three generic conics $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ in the complex projective plane $\mathbb{P}^{2}$, where $X_{0}=X:=\mathbb{P}^{2}$, $D_{0}:=\mathcal{C}=\sum_{i=1}^{3}\mathcal{C}_{i}$ is a simple normal crossing divisors on $X$ whose irreducible components $\mathcal{C}_{i}$ are smooth conics and $V_{0}:=T_{X}(-\mathrm{log}\,D_{0})$. Define $X_{1}:=\mathbb{P}(V_{0})$ with the natural projection $\pi_{1,0}:X_{1}\rightarrow X_{0}$. Set $D_{1}:=\pi_{1,0}^{*}D_{0}$. The subbundle $V_{1}\subset T_{X_{1}}(-\mathrm{log}\,D_{1})$ is defined fiberwise by:
(2.4) $V_{1,(x,[v])}:=\{\xi\in T_{X_{1},(x,[v])}(-\mathrm{log}\,D_{1})\,|\,\left(\pi_{1,0}\right)_{*}\xi\in\mathbb{C}\cdot v\},$
with $\mathbb{C}\cdot v\subset V_{0,x}=T_{X_{0},x}(-\mathrm{log}\,D_{0})$. Thus we obtain the first level $(X_{1},D_{1},V_{1})$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C})$.
Proceeding likewise, we can construct the second level of logarithmic Demailly–Semple tower:
$(X_{2},D_{2},V_{2})\rightarrow(X_{1},D_{1},V_{1})\rightarrow(X_{0},D_{0},V_{0}),$
together with the projections $\pi_{2,1}\colon X_{2}\rightarrow X_{1}$ and $\pi_{2,0}\colon X_{2}\rightarrow X_{0}$.
Let $\mathcal{O}_{X_{1}}(-1)$ denote the tautological line bundle of $X_{1}=\mathbb{P}(V_{0})$ such that $\mathcal{O}_{X_{1}}(-1)_{(x,[v])}=\mathbb{C}\cdot v$. By definition, there is a short exact sequence
[…]
where $T_{X_{1}/X_{0}}$ is the relative tangent bundle of the fibration $\pi_{1,0}\colon X_{1}\rightarrow X_{0}$. Moreover, there is a relative version of the Euler sequence of the tangent bundle of the fiber $\mathbb{P}(V_{0,x})$:
[…]
Furthermore, there exists a canonical injection $\mathcal{O}_{X_{2}}(-1)\hookrightarrow\pi_{2,1}^{*}V_{1}$. Combining this with the above sequence yields a natural morphism of line bundles:
[…]
whose zero divisor is the hyperplane subbundle:
(2.7) $\Gamma_{2}\coloneqq\mathbb{P}(T_{X_{1}/X_{0}})\subset\mathbb{P}(V_{1})=X_{2}.$
Consequently, we can obtain the isomorphism:
(2.8) $\mathcal{O}_{X_{2}}(\Gamma_{2})\cong\pi_{2,1}^{*}\mathcal{O}_{X_{1}}(-1)\otimes\mathcal{O}_{X_{2}}(1).$
Let $f\colon\mathbb{D}_{R}\rightarrow X_{0}$ be a nonconstant holomorphic map from the disc $\mathbb{D}_{R}=\{z\in\mathbb{C}:|z|<R\}$ of radius $R\in\mathbb{R}_{+}\cup\{\infty\}$. To define its liftings to the first level $X_{1}$ of the logarithmic Demailly–Semple tower, we consider the following two cases for $f^{\prime}(t)\in T_{X_{0},f(t)}$:
When $0\neq f^{\prime}(t)\in T_{X_{0},f(t)}\left(-\mathrm{log}\,D_{0}\right)$, the point $f^{\prime}(t)]\in\mathbb{P}\left(T_{X_{0},f(t)}\left(-\mathrm{log}\,D_{0}\right)\right)$ is well defined;
When $f^{\prime}(t_{0})=0$ or $f^{\prime}(t_{0})\in T_{X_{0},f(t_{0})}\setminus T_{X_{0},f(t_{0})}\left(-\mathrm{log}\,D_{0}\right)$, we can define $f^{\prime}(t_{0})]\in\mathbb{P}\left(T_{X_{0},f(t_{0})}\left(-\mathrm{log}\,D_{0}\right)\right)$ as following: near $t_{0}\in\mathbb{D}_{R}$, choose a local frame $(e_{1},e_{2})$ for $T_{X_{0}}(-\mathrm{log}\,D_{0})$ in a neighborhood of $f(t_{0})$. In this frame, the derivative $f^{\prime}(t)$ has a local expression $f^{\prime}(t)=u_{1}(t)e_{1}+u_{2}(t)e_{2}$, where $u_{1}(t),u_{2}(t)$ are meromorphic functions with at most finite-order poles. Let $\ell\in\mathbb{Z}$ be the minimum vanishing order (equivalently, the negative of the maximal pole order) of $u_{1}(t)$ and $u_{2}(t)$ at $t_{0}$. Factoring out $(t-t_{0})^{\ell}$, we can write
$f^{\prime}(t)=(t-t_{0})^{\ell}\bigl(v_{1}(t)e_{1}+v_{2}(t)e_{2}\bigr),$
where $v_{1}(t)$ and $v_{2}(t)$ are holomorphic near $t_{0}$ and not simultaneously vanishing at $t_{0}$. We define the projective class:
$f^{\prime}(t_{0})]\coloneqq[v_{1}(t_{0}):v_{2}(t_{0})]\in\mathbb{P}(T_{X_{0},f(t_{0})}(-\mathrm{log}\,D_{0})).$
This definition is independent of the choice of local frame, so $f^{\prime}(t_{0})]\in\mathbb{P}\left(T_{X_{0},f(t_{0})}\left(-\mathrm{log}\,D_{0}\right)\right)$ is well defined.
With this convention, the first lifting $f_{[1]}=(f;[f^{\prime}])$ of $f$ is a well-defined holomorphic map from $\mathbb{D}_{R}$ to $X_{1}$. Repeating the process for $f_{[1]}$, the second lifting $f_{[2]}=(f_{[1]};[f_{[1]}^{\prime}])$ of $f$ is also well-defined and holomorphic from $\mathbb{D}_{R}$ to $X_{2}$.
Moreover, from the definition of $\Gamma_{2}$ and the identity $\pi_{1,0}\circ f_{[1]}=f$, it follows that $f_{[2]}(t)\in\Gamma_{2}$ if and only if $(\pi_{1,0})_{*}f_{[1]}^{\prime}(t)=f^{\prime}(t)=0$.
Demailly [Demailly1997, Corollary 5.12] observed that, for any $w_{0}\in X_{2}$, there exists an open neighborhood $U_{w_{0}}\subset X_{2}$ such that for all the $w\in U_{w_{0}}$, one can find a holomorphic family of germs $f_{w}\colon(\mathbb{C},0)\rightarrow U$ such that $\left(f_{w}\right)_{[2]}(0)=w$ and $\left(f_{w}\right)_{[1]}^{\prime}(0)\neq 0$. By the construction of the Demailly–Semple tower, we have $f_{w}^{\prime}(t)\in\mathcal{O}_{X_{1}}(-1)_{\left(f_{w}\right)_{[1]}(t)}$ and $\left(f_{w}\right)_{[1]}^{\prime}(t)\in\mathcal{O}_{X_{2}}(-1)_{\left(f_{w}\right)_{[2]}(t)}$. Then for any $P\in E_{2,m}T_{X}^{*}(\mathrm{log}\,D)_{x}$ with $x\in\pi_{2,0}(U_{w_{0}})\subset X_{0}$, this induces a holomorphic section $\sigma_{P}(w)$ of $\mathcal{O}_{X_{2}}(m)$ over the fiber $X_{2,x}$ defined by:
$\sigma_{P}(w)\coloneqq P(f_{w}^{\prime},f_{w}^{\prime\prime})(0)\left(\left(f_{w}\right)_{[1]}^{\prime}(0)\right)^{-m}.$
More generally, the following fundamental isomorphism holds:
Direct image formula (cf. [Demailly1997, Theorm 6.8], [Dethloff-Lu2001, Propsition 3.9]).
The direct image sheaf
(2.9) $\left(\pi_{2,0}\right)_{*}\mathcal{O}_{X_{2}}(m)\cong\mathcal{O}\left(E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\right)$
is isomorphic to the sheaf of invariant logarithmic $2$-jet differentials of weighted degree $m$.
2.3. A Bogomolov-type Vanishing Theorem
In the subsequent Sections 3 and 4, the existence of negatively twisted logarithmic $1$-jet and $2$-jet differentials will be established via Riemann–Roch calculations. This reduces to verifying the vanishing of the corresponding second cohomology groups:
$H^{2}\big(\mathbb{P}^{2},S^{m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)\quad\text{and}\quad H^{2}\big(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big),$
respectively. A key tool in proving these vanishing results is the following Bogomolov-type vanishing theorem.
Theorem 2.1 (cf. [Chen2025, Theorem 4.4]).
Let $(X,D)$ be a log pair where $X$ is a smooth projective surface and $D$ is a reduced simple normal crossing divisor. Assume that $(X,D)$ is of general type (i.e., $K_{X}+D$ is big) and $K_{X}+D$ is nef. Then
$H^{0}\big(X,S^{p}T_{X}(-\mathrm{log}\,D)\otimes\mathcal{O}_{X}(q(K_{X}+D))\big)=0,$
for all integers $p,q$ satisfying $p>2q$.
In the compact case, Bogomolov [Bogomolov1978] established a seminal vanishing theorem under the assumption that tangent bundle $T_{X}$ is semi-stable. Later, Enoki [Enoki1987] showed that $T_{X}$ is $K_{X}$-semi-stable for minimal Kähler space $X$ with big $K_{X}$, and Tsuji [Tsuji1988] proved that $T_{X}$ is $K_{X}$-semi-stable for smooth minimal algebraic variety (i.e., $K_{X}$ is nef). Combining these results, the Bogomolov-type vanishing theorem holds for projective variety $X$ with big and nef $K_{X}$; see [Demailly1997, Theorem 14.1].
3. Proof of Theorem 1.3
Let $\mathcal{C}=\sum_{i=1}^{3}\mathcal{C}_{i}$ be a simple normal crossing divisor of total degree $d=d_{1}+d_{2}+d_{3}$, where $d_{i}=\deg\mathcal{C}_{i}$ satisfy $d_{1}\geqslant d_{2}\geqslant d_{3}\geqslant 2$ and $d_{1}\geqslant 3.$ To derive an effective Second Main Theorem, we seek nontrivial global sections of the twisted sheaf
$S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(1)^{-(d-3)\delta m}\cong S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\,-\delta m},$
where $\delta>0$ is a small rational number to be chosen, and $m$ is taken sufficiently divisible.
Definition 3.1 (cf. [Demailly-Elgoul2000, p. 517], [Elgoul2003, Definition 1.3.1]).
The $1$-jet log threshold of $(\mathbb{P}^{2},\mathcal{C})$ is defined by
$\overline{\theta}_{1}(\mathbb{P}^{2},\mathcal{C})\coloneqq\sup\left\{\frac{t}{m}\,\middle|\,m,t>0,\;H^{0}\big(\mathbb{P}^{2},S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)\neq 0\right\}.$
If no such pair $(m,t)$ exists, the threshold is defined to be zero.
Proposition 3.2 (Proof of (1.4)).
Let $\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i}$ be a union of three smooth curves $\mathcal{C}_{i}$ of degrees $d_{i}$ with simple normal crossings. Assume that $d_{1}\geqslant d_{2}\geqslant d_{3}\geqslant 2$ and $d_{1}\geqslant 3$. Then one has the estimate
$0<\frac{1}{\overline{\theta}_{1}(\mathbb{P}^{2},\mathcal{C})}\leqslant\frac{1}{(d-3)\delta_{1}},$
where
(3.1) $\delta_{1}=\frac{1}{2}-\dfrac{\sqrt{9(d-1)^{2}-12(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})}}{6(d-3)}>0.$
Proof.
We follow the standard Riemann-Roch argument (cf. [Demailly-Elgoul2000, Elgoul2003]).
By Serre duality, we obtain
$H^{2}\big(\mathbb{P}^{2},S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\big)$ $=H^{0}\big(\mathbb{P}^{2},K_{\mathbb{P}^{2}}\otimes S^{m}T_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m}\big)$
$=H^{0}\big(\mathbb{P}^{2},S^{m}T_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m-\frac{3}{(d-3)}}\big).$
For any rational number $\delta\in(0,\frac{1}{2})$ and sufficiently large $m$, the inequality $m>2(\delta m-\frac{3}{d-3})$ holds. Applying the Bogomolov-type Vanishing Theorem 2.1, we obtain
$H^{2}\big(\mathbb{P}^{2},S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\big)=0$
which yields
$\dim H^{0}\Big(\mathbb{P}^{2},S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\Big)\geqslant\chi\Big(S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\Big),$
for large $m$. A standard Riemann-Roch computation gives
$\chi\Big(S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\Big)$ $=\dfrac{1}{6}\big((m-\delta m)^{3}-(-\delta m)^{3}\big)\bar{c}_{1}^{2}-\dfrac{1}{6}m^{3}\bar{c}_{2}+O(m^{2})$
$=\dfrac{1}{2}m^{3}\bigg((d-3)^{2}\delta^{2}-(d-3)^{2}\delta+\frac{1}{3}(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})-d+2\bigg)+O(m^{2}).$
The leading coefficient $\varphi(\delta)=(d-3)^{2}\delta^{2}-(d-3)^{2}\delta+\frac{1}{3}(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})-d+2$ has two positive roots given by
$\delta_{1}=\frac{1}{2}-\dfrac{\sqrt{9(d-1)^{2}-12(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})}}{6(d-3)},\quad\quad\delta_{2}=\frac{1}{2}+\dfrac{\sqrt{9(d-1)^{2}-12(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})}}{6(d-3)}>0.$
We can check the positivity of $\delta_{1}$ directly. Hence $\varphi(\delta)>0$ for all $\delta\in(0,\delta_{1})$. Thus we conclude the proof. ∎
4. Twisted Invariant Logarithmic $2$-Jet Differentials and Threshold
To obtain a constant $c$ as small as possible in the Second Main Theorem for three conics $\mathcal{C}\subset\mathbb{P}^{2}$, by Theorem (1.2), we seek negatively twisted invariant logarithmic $2$-jet differentials with relatively high vanishing order $t$ compared to the degree $m$. This motivates the study of the following threshold:
Definition 4.1 (cf. [Demailly-Elgoul2000, p. 517], [Elgoul2003, Definition 1.3.1]).
The $2$-jet log threshold of $(\mathbb{P}^{2},\mathcal{C})$ is defined by
$\overline{\theta}_{2}(\mathbb{P}^{2},\mathcal{C})\coloneqq\sup\left\{\frac{t}{m}\,\middle|\,m,t>0,\;H^{0}\big(X_{2},\mathcal{O}_{X_{2}}(m)\otimes\pi_{2,0}^{*}\mathcal{O}_{X}(-t)\big)\neq 0\right\}.$
If no such pair $(m,t)$ exists, the threshold is defined to be zero.
If $\overline{\theta}_{2}>0$, then for any sufficiently small $\epsilon>0$, there exists a negatively twisted logarithmic $2$-jet differential
$\omega\in H^{0}\big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)$
with ratio $\overline{\theta}_{2}-\epsilon<\frac{t}{m}\leqslant\overline{\theta}_{2}$. Furthermore, if $f^{*}\omega\not\equiv 0$ also holds, then Theorem 1.2 yields the Second Main Theorem with constant
$\frac{1}{\overline{\theta}_{2}}\leqslant c=\frac{m}{t}<\frac{1}{\overline{\theta}_{2}-\epsilon}.$
The significance of the $2$-jet log threshold stems from its role as the fundamental barrier that dictates the possibility of constructing jet differentials with a prescribed ratio of vanishing orders to weighted degrees.
The following result establishes an explicit upper bound for $\frac{1}{\overline{\theta}_{2}}$ in the case of three smooth conics with simple normal crossings:
Proposition 4.2.
Let $\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i}$ be a union of three smooth conics $\mathcal{C}_{i}$ with simple normal crossings. Then one has the estimate
$0<\frac{1}{\overline{\theta}_{2}(\mathbb{P}^{2},\mathcal{C})}<\frac{3}{4-\sqrt{10}}=3.58\ldots.$
The proof employs a Riemann–Roch calculation, which is standard (cf. [Demailly-Elgoul2000]).
Proof.
We seek nonzero invariant logarithmic $2$-jet differential forms with high vanishing order, which correspond to global sections of the twisted vector bundle
$E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(1)^{-3\delta m}\cong E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m},$
where the logarithmic canonical line bundle
$\overline{K}_{\mathbb{P}^{2}}=K_{\mathbb{P}^{2}}\otimes\mathcal{O}(\mathcal{C})\cong\mathcal{O}_{\mathbb{P}^{2}}(-3)\otimes\mathcal{O}_{\mathbb{P}^{2}}(6)\cong\mathcal{O}_{\mathbb{P}^{2}}(3),$
with $\delta>0$ a rational number to be chosen and $m$ sufficiently divisible.
The standard approach begins with Serre duality:
(4.1) $H^{2}\big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\big)=H^{0}\big(\mathbb{P}^{2},K_{\mathbb{P}^{2}}\otimes E_{2,m}T_{\mathbb{P}^{2}}(-\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m}\big).$
The bundle $K_{\mathbb{P}^{2}}\otimes E_{2,m}T_{\mathbb{P}^{2}}(-\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m}$ admits a filtration, which is induced by the dual of (2.3) with a twist, with graded pieces
$S^{m-3j}T_{\mathbb{P}^{2}}(-\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-j}\otimes K_{\mathbb{P}^{2}}\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m}=S^{m-3j}T_{\mathbb{P}^{2}}(-\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m-j-1}$
for $0\leqslant j\leqslant\frac{m}{3}$.
For any rational number $\delta\in(0,\frac{1}{3})$ and sufficiently divisible $m\gg 1$, the inequality
$m-3j>2(\delta m-j-1)$
automatically holds for all $0\leqslant j\leqslant\frac{m}{3}$. By applying the Bogomolov-type Vanishing Theorem 2.1, we deduce:
$H^{0}\Big(\mathbb{P}^{2},S^{m-3j}T_{\mathbb{P}^{2}}(-\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\delta m-j-1}\Big)=0\quad\,\,\,\text{for all }\quad 0\leqslant j\leqslant\tfrac{m}{3}.$
This establishes the vanishing of the right-hand side in (4.1), which consequently implies the vanishing of the left-hand side as well. We thus obtain the lower bound
$\dim H^{0}\Big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\Big)\geqslant\chi\Big(E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\Big),$
since the Euler characteristic decomposes as
$\chi(\cdots)=\dim H^{0}(\cdots)-\dim H^{1}(\cdots)+\dim H^{2}(\cdots)=\dim H^{0}(\cdots)-\dim H^{1}(\cdots).$
To estimate the asymptotic growth of the Euler characteristic, we employ the Schur bundle filtration (see (2.3)):
[…]
(4.2) $=\bigoplus_{0\leqslant b\leqslant\frac{m}{3}}\Gamma^{(m-2b,\,b)}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}).$
Together with the Pieri formula (cf. [FultonHarris1991, p. 455], [Demailly1997, p. 66]):
$\Gamma^{(\ell_{1},\,\ell_{2})}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}=\Gamma^{(\ell_{1}-\delta m,\,\ell_{2}-\delta m)}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}),$
we derive the following computation (cf. [Merker2015, p. 77]):
$\chi\big(E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{-\delta m}\big)$
$=\,$ $\sum_{b=0}^{\frac{m}{3}}\chi\big(\Gamma^{(m-2b-\delta m,\,b-\delta m)}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\big)$
$=\,$ $\bigg(\frac{1}{6}\int_{0}^{\frac{m}{3}}\big[(m-2b-\delta m)^{3}-(b-\delta m)^{3}\big]\mathrm{d}b\bigg)\bar{c}_{1}^{2}\,-\,\bigg(\frac{1}{6}\int_{0}^{\frac{m}{3}}(m-3b)^{3}\mathrm{d}b\bigg)\bar{c}_{2}+O(m^{3})$
$=\,$ $\frac{m^{2}(m+3)}{108}(54\delta^{2}m-48\delta m+13m-18\delta+15)\bar{c}_{1}^{2}-\frac{m^{4}}{72}\bar{c}_{2}+O(m^{3})$
$=\,$ $\frac{m^{4}}{648}\big((54\delta^{2}-48\delta+13)\bar{c}_{1}^{2}-9\bar{c}_{2}\big)+O(m^{3}).$
Here the logarithmic Chern numbers are determined by the formula $\bar{c}_{1}^{2}=(d-3)^{2}=9$ and $\bar{c}_{2}=d^{2}-3d+3-\sum_{i\neq j}d_{i}d_{j}=9$, where $d_{1}=d_{2}=d_{3}=2$ and $d=d_{1}+d_{2}+d_{3}=6$.
The quadratic leading coefficient (after removing the common factor $\frac{9}{648}$)
$\varphi(\delta):=(54\delta^{2}-48\delta+13)-9=54\delta^{2}-48\delta+4$
has two positive roots given by $\delta_{1}=\frac{4-\sqrt{10}}{9}$ and $\delta_{2}=\frac{4+\sqrt{10}}{9}$.
Hence, $\varphi(\delta)>0$ for all $\delta\in(0,\delta_{1})$. It follows that $\overline{\theta}_{2}\geqslant 3\delta_{1}=0.279\ldots$, which completes the proof of Proposition 4.2. ∎
To apply Siu’s strategy of slanted vector fields, we follow the argument of Siu [Siu2004, pp. 558–559] to extend a negatively twisted invariant logarithmic $2$-jet differential as a holomorphic family, invoking the following standard result in algebraic geometry:
Theorem 4.3 ([HartshorneGTM52, p. 288]).
Let $\tau\colon\mathcal{Z}\rightarrow S$ be a flat holomorphic family of compact complex spaces and let $\mathcal{L}\rightarrow\mathcal{Z}$ be a holomorphic vector bundle. Then there exists a proper subvariety $Z\subset S$ such that for each $s_{0}\in S\backslash Z$, the restriction map $H^{0}\big(\tau^{-1}\left(U_{s_{0}}\right),\mathcal{L}\big)\rightarrow H^{0}\big(\tau^{-1}(s_{0}),\mathcal{L}|_{\tau^{-1}(s_{0})}\big)$ is onto, for some Zariski-dense open set $U_{s_{0}}\subset S$ containing $s_{0}$.
Now we apply Demailly–El Goul’s argument (cf. [Demailly-Elgoul2000, p. 532], [Elgoul2003, Lemma 1.3.2]) to show the existence of “irreducible” negatively twisted invariant logarithmic $2$-jet differentials.
Proposition 4.4.
For some positive integer $m,t\geqslant 1$ with m<5t{m}<{5}t, there exists a Zariski-open subset $U^{m,t}$ in the parameter space $\mathbb{P}^{5}\times\mathbb{P}^{5}\times\mathbb{P}^{5}$ (see (5.1)) of three conics $\mathcal{C}$ in $\mathbb{P}^{2}$, such that:
  1. (1)
    For each parameter $a\in U^{m,t}$, the corresponding $\mathcal{C}_{a}$ forms a simple normal crossing divisor.
  2. (2)
    There is a holomorphic family
    $\left\{\omega_{1,a}\;\middle|\;0\not\equiv\omega_{1,a}\in H^{0}\left(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}_{a})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\right),\;a\in U^{m,t}\right\},$
    which varies holomorphically with respect to the variable $a\in U^{m,t}$. Moreover, each zero locus $\{\omega_{1,a}=0\}$ on the second level $X_{2,a}$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C}_{a})$ is irreducible and reduced modulo $\Gamma_{2,a}$ (see (2.7)).
Proof.
First, for any integers $M,T\geqslant 1$, Theorem 4.3 guarantees that there exists a proper Zariski-closed subset $Z_{M,T}$ in the parameter space $\mathbb{P}^{5}\times\mathbb{P}^{5}\times\mathbb{P}^{5}$ of three conics, such that, for any parameter $a$ outside $Z_{M,T}$, every nonzero global section of $E_{2,M}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}_{a})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-T)$ extends to a holomorphic family parameterized by some Zariski open neighborhood of $a$. Thus, to ensure the local extendability of any nonzero negatively twisted logarithmic $2$-jet differential on $(\mathbb{P}^{2},\mathcal{C}_{a_{0}})$, we choose a point $a_{0}$ in $(\mathbb{P}^{5}\times\mathbb{P}^{5}\times\mathbb{P}^{5})\setminus\cup_{M,T\geqslant 1}Z_{M,T}$ such that $\mathcal{C}_{a_{0}}$ is a simple normal crossing divisor. The existence of such $a_{0}$ is guaranteed by the Bertini theorem.
By Proposition 4.2, noting that $5>\frac{3}{4-\sqrt{10}}=3.58\ldots$, for some integer $M,T\geqslant 1$ with ratio $\frac{1}{5}<\frac{T}{M}\leqslant\overline{\theta}_{2}(\mathbb{P}^{2},\mathcal{C}_{a_{0}})$, there exists a nonzero negatively twisted invariant logarithmic $2$-jet differential
$\omega_{0}\in H^{0}\big(\mathbb{P}^{2},E_{2,M}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}_{a_{0}})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-T)\big).$
Next, we decompose the zero locus $\{\omega_{0}=0\}$ in the second level $X_{2,a_{0}}$ of the logarithmic Demailly-Semple tower of $(\mathbb{P}^{2},\mathcal{C}_{a_{0}})$ into irreducible components. By the same argument as Demailly and El Goul [Demailly-Elgoul2000, p. 532], [Elgoul2003, Lemma 1.3.2], we can extract an irreducible and reduced component $Z$, which in fact “corresponds” to some irreducible negatively twisted invariant logarithmic $2$-jet differential
(4.3) $\omega_{1}\in H^{0}\big(\mathbb{P}^{2},E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}_{a_{0}})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big),\quad\text{with}\ \tfrac{t}{m}\geqslant\tfrac{T}{M}>\tfrac{1}{5},$
in the sense that $\{\omega_{1}=0\}=Z+b\Gamma_{2,a_{0}}$ (see (2.7) and (2.8)) for some integer $b>0$. The argument is as follows.
Let $u_{1}=\pi_{2,1}^{*}\mathcal{O}_{X_{1,a_{0}}}(1)$, $u_{2}=\mathcal{O}_{X_{2,a_{0}}}(1)$ and $h=\mathcal{O}_{\mathbb{P}^{2}}(1)$. Then the zero divisor $Z_{\omega_{0}}=\{\omega_{0}=0\}\subset X_{2,a_{0}}$ is linearly equivalent to $Mu_{2}-T\pi_{2,0}^{*}h$ in the Picard group
$\operatorname{Pic}(X_{2,a_{0}})\cong\operatorname{Pic}(\mathbb{P}^{2})\oplus\mathbb{Z}u_{1}\oplus\mathbb{Z}u_{2}\cong\mathbb{Z}^{3}.$
Let $Z_{\omega_{0}}=\sum p_{j}Z_{j}$ be the irreducible decomposition of the divisor $Z_{\omega_{0}}$, with each $Z_{j}$ irreducible and reduced. Decompose each component $Z_{j}$ linear equivalently in the Picard group $\operatorname{Pic}(X_{2,a_{0}})$ as
$Z_{j}\sim b_{1,j}u_{1}+b_{2,j}u_{2}-t_{j}\pi_{2,0}^{*}h,\qquad b_{1,j}\;,\;b_{2,j}\;,\;t_{j}\in\mathbb{Z}.$
By [Demailly-Elgoul2000, Lemma 3.3], the effectiveness of $Z_{j}$ implies that exactly one of the following mutually exclusive cases holds:
  • $\left(b_{1,j},b_{2,j}\right)=(0,0)$ and $Z_{j}\in\pi_{2,0}^{*}\operatorname{Pic}\left(\mathbb{P}^{2}\right)$, $t_{j}>0$;
  • $\left(b_{1,j},b_{2,j}\right)=(-1,1)$, then $Z_{j}$ contains $\Gamma_{2,,a_{0}}$, where $\Gamma_{2,,a_{0}}$ is an irreducible divisor on $X_{2,a_{0}}$. Thus $Z_{j}=\Gamma_{2,a_{0}}$ and $t_{j}=0$;
  • $b_{1,j}\geqslant 2b_{2,j}\geqslant 0$ and $m_{j}\coloneqq b_{1,j}+b_{2,j}>0$.
In the third case, the irreducible and reduced divisor $Z_{j}$ corresponds a section
$\omega_{j}\in H^{0}\big(X_{2,a_{0}},\pi_{2,1}^{*}\mathcal{O}_{X_{1,a_{0}}}(b_{1,j})\otimes\mathcal{O}_{X_{2,a_{0}}}(b_{2,j})\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-t_{j})\big).$
Noting that $\pi_{2,1}^{*}\mathcal{O}_{X_{1,a_{0}}}(b_{1})\otimes\mathcal{O}_{X_{2,a_{0}}}(b_{2})\cong\mathcal{O}_{X_{2,a_{0}}}\left(b_{1}+b_{2}\right)\otimes\mathcal{O}_{X_{2,a_{0}}}\left(-b_{1}\Gamma_{2,a_{0}}\right)$ (see (2.8)), we may regard
$\omega_{j}\in H^{0}\big(X_{2,a_{0}},\mathcal{O}_{X_{2,a_{0}}}(m_{j})\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-t_{j})\big),$
whose zero divisor is $Z_{j}+b_{1,j}\Gamma_{2,a_{0}}$. Since both $Z_{j}$ and $\Gamma_{2,a_{0}}$ are irreducible and reduced, $\omega_{j}$ can not be factorized into a product of other logarithmic $2$-jet differentials.
Let $\max\{{t_{j}}/{m_{j}}\}$ denote the maximum ratio among those sections obtained from the third case. Since $M=\sum p_{j}m_{j}$ and $T=\sum p_{j}t_{j}$, we have $\max\{{t_{j}}/{m_{j}}\}\geqslant{T}/{M}>1/5$. Without loss of generality, we may assume that $t_{1}/m_{1}=\max\{t_{j}/m_{j}\}$, and denote $m\coloneqq m_{1}$ and $t\coloneqq t_{1}$. By applying the direct image formula (2.9) and abusing the notation for $\omega_{1}$, we thus obtain (4.3).
By our choice of $a_{0}\notin\cup_{M,T\geqslant 1}Z_{M,T}$, automatically $\omega_{1}$ extends to a holomorphic family
$\left\{\omega_{1,a}\;\middle|\;0\not\equiv\omega_{1,a}\in H^{0}\left(\mathbb{P}^{2},E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}_{a})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\right),\;a\in U_{a_{0}}^{m,t}\right\},$
which varies holomorphically with respect to the parameter $a$ in some Zariski-open neighborhood $U_{a_{0}}^{m,t}$ of $a_{0}$, with $\omega_{1,a_{0}}=\omega_{1}$. By shrinking to a smaller Zariski-open subset $U^{m,t}\subset U_{a_{0}}^{m,t}$ if necessary, we may ensure that each $a\in U^{m,t}$ defines three conics $\mathcal{C}_{a}$ with simple normal crossings while the zero locus $\{\omega_{1,a}=0\}$ on the second level $X_{2,a}$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C}_{a})$ is irreducible and reduced modulo $\Gamma_{2,a}$. ∎
5. Generating Independent Twisted Jet Differentials via Slanted Vector Fields
To construct an additional family of independent negatively twisted jet differentials, we employ Siu’s method of slanted vector fields [Siu2004]. This approach generalizes to $k$-jet spaces the original technique developed for $1$-jet spaces by Clemens [Clemens1986], Ein [Ein1988], and Voisin [Voisin1996] in their study of rational curves.
The method was first implemented by Păun [mihaipaun2008] for compact surfaces in $\mathbb{P}^{3}$. Rousseau [Rousseau2009] subsequently established the logarithmic theory for pairs $(\mathbb{P}^{2},D)$ with curve configurations $D$, while in [Rousseau2007cpt, Rousseau2007log] he further extended both Păun’s setting and the logarithmic case to threefolds. For hypersurfaces $H\subset\mathbb{P}^{n}$ of arbitrary dimension, Merker [Merker2009] established the complete generalization, thereby realizing Siu’s original vision concerning generic global generation. The logarithmic counterpart for pairs $(\mathbb{P}^{n},H)$ was subsequently obtained by Darondeau [Darondeau_vectorfield].
This section adapts the standard technique of Siu’s slanted vector fields to the case of three generic conics in $\mathbb{P}^{2}$. The content is entirely classical and no originality is claimed.
5.1. Logarithmic Slanted Vector Fields for Three Conics in $\mathbb{P}^{2}$
The framework presented in this subsection specializes Rousseau’s logarithmic setting in $\mathbb{P}^{2}$ to configurations of three conics. We emphasize that this subsection contains no original contributions and is included solely for completeness and the reader’s convenience.
Let
(5.1) $\mathcal{X}=\cup_{\lambda=1}^{3}\mathcal{X}_{\lambda}\subset\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$
denote the universal family of algebraic curves in $\mathbb{P}^{2}$ consisting of three components of degrees $d_{1}$, $d_{2}$, and $d_{3}$, respectively. Here, each $\mathcal{X}_{\lambda}$ is the universal curve of degree $d_{\lambda}$, defined by the equation $\sum_{|\alpha|=d_{\lambda}}A_{\alpha}^{\lambda}Z^{\alpha}=0$, where $A^{\lambda}]=[A_{\alpha}^{\lambda}]_{|\alpha|=d_{\lambda}}\in\mathbb{P}^{N_{d_{\lambda}}}$ are the homogeneous coefficients and $Z]=[Z_{0}:Z_{1}:Z_{2}]\in\mathbb{P}^{2}$ are the projective coordinates. For a multi-index $\alpha=(\alpha_{0},\alpha_{1},\alpha_{2})\in\mathbb{N}^{3}$ with length $|\alpha|:=\alpha_{0}+\alpha_{1}+\alpha_{2}$, we employ the notation $Z^{\alpha}:=Z_{0}^{\alpha_{0}}Z_{1}^{\alpha_{1}}Z_{2}^{\alpha_{2}}$. The dimension of the parameter space for each component is given by $N_{d_{\lambda}}=\binom{d_{\lambda}+2}{2}-1$.
For any integer $k\geqslant 1$, the vertical logarithmic $k$-jet space is then defined as
$\overline{J_{k}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right):=\ker\left(\mathrm{pr}_{2}\right)_{*}\subset J_{k}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}},-\mathrm{log}\,\mathcal{X}\right),$
where $\mathrm{pr}_{2}\colon\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\longrightarrow{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$ denotes the natural projection onto the parameter space and $J_{k}\big(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}},-\mathrm{log}\,\mathcal{X}\big)$ is the logarithmic $k$-jet bundle along the universal family $\mathcal{X}$.
We define the universal intersection family
$\mathcal{Y}=\cap_{\lambda=1}^{3}\mathcal{Y}_{\lambda}\subset\mathbb{P}^{5}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}+1},$
where for each $1\leqslant\lambda\leqslant 3$, the hypersurface $\mathcal{Y}_{\lambda}$ is the universal family of degree $d_{\lambda}$ hypersurfaces in $\mathbb{P}^{5}$, given explicitly by $\mathcal{Y}_{\lambda}:=\left\{A_{\mathbf{0}}^{\lambda}Z_{2+\lambda}^{d_{\lambda}}+\sum_{|\alpha|=d_{\lambda}}A_{\alpha}^{\lambda}Z^{\alpha}=0\right\}\subset\mathbb{P}^{5}\times\mathbb{P}^{N_{d_{\lambda}}+1}$. Here, the space $\mathbb{P}^{5}$ has homogeneous coordinates $Z:Z_{3}:Z_{4}:Z_{5}]=[Z_{0}:\cdots:Z_{5}]$, and the parameter space $\mathbb{P}^{N_{d_{\lambda}}+1}$ is equipped with homogeneous coordinates $A^{\lambda}:A_{\mathbf{0}}^{\lambda}]=[\ldots:A_{\alpha}^{\lambda}:\ldots:A_{\mathbf{0}}^{\lambda}]_{|\alpha|=d_{\lambda}}$.
For notational convenience, we denote the exceptional locus $\{Z=\mathbf{0}\}\cup\cup_{\lambda=1}^{3}\{A^{\lambda}=\mathbf{0}\}$, where $\{Z=\mathbf{0}\}$ denotes the subset of $\mathbb{P}^{5}$ defined by $Z_{0}=Z_{1}=Z_{2}=0$ and $\{A^{\lambda}=\mathbf{0}\}$ refers to the locus in $\mathbb{P}^{N_{d_{\lambda}}+1}$ where all coordinates vanish except $A_{\mathbf{0}}^{\lambda}$. There exists a natural forgetful map
$\pi\colon\mathbb{P}^{5}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}+1}\setminus\left(\{Z=\mathbf{0}\}\cup\cup_{\lambda=1}^{3}\{A^{\lambda}=\mathbf{0}\}\right)\longrightarrow\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$
that forgets the auxiliary coordinates. Explicitly, $\pi$ is given by
$\left([Z:Z_{3}:Z_{4}:Z_{5}],[A^{1}:A_{\mathbf{0}}^{1}],[A^{2}:A_{\mathbf{0}}^{2}],[A^{3}:A_{\mathbf{0}}^{3}]\right)\longmapsto\left([Z],[A^{1}],[A^{2}],[A^{3}]\right).$
We observe the containment relation
(5.2) $\mathcal{Y}\cap\left(\{Z=\mathbf{0}\}\cup\cup_{\lambda=1}^{3}\{A^{\lambda}=\mathbf{0}\}\right)\subset\mathcal{Y}\cap\left(\cup_{\lambda=1}^{3}\left(\{A_{\mathbf{0}}^{\lambda}=0\}\cup\{A^{\lambda}=\mathbf{0}\}\right)\right),$
from the inclusion $\mathcal{Y}\cap\{Z=\mathbf{0}\}\subset\mathcal{Y}\cap\cup_{\lambda=1}^{3}\{A_{\mathbf{0}}^{\lambda}=0\}$, which is deduced by the definition of $\mathcal{Y}$.
Now we consider $\mathcal{Y}^{*}\coloneqq\mathcal{Y}\setminus\left(\cup_{\lambda=1}^{3}\left(\{A_{\mathbf{0}}^{\lambda}=0\}\cup\{A^{\lambda}=\mathbf{0}\}\right)\right)$ on which the restricted forgetful map
$\pi|_{\mathcal{Y}^{*}}\colon\mathcal{Y}^{*}\rightarrow\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$
is well-defined according to the containment relation (5.2). Furthermore, we obtain the identification of the preimage $\left(\pi|_{\mathcal{Y}^{*}}\right)^{-1}(\mathcal{X})=\{Z_{3}Z_{4}Z_{5}=0\}=:\mathcal{D}$. This shows that $\pi$ induces a logarithmic morphism from $\left(\mathcal{Y}^{*},\mathcal{D}\right)$ to $\big(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}},\mathcal{X}\big)$, which consequently gives rise to a dominant map between jet spaces:
(5.3) $\pi_{[2]}\colon\overline{J_{2}^{v}}(\mathcal{Y}^{*})\longrightarrow\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right).$
Here, $\overline{J_{2}^{v}}(\mathcal{Y}^{*})$ represents the space of vertical logarithmic 2-jets with respect to the normal crossing divisor $\mathcal{D}=\{Z_{3}Z_{4}Z_{5}=0\}$. This construction allows us to perform all local computations on the more tractable space $\mathcal{Y}^{*}$ through pullback via $\pi$, where we have access to explicit coordinate expressions.
On the affine chart $\{Z_{0}\neq 0\}\times{\textstyle\prod}_{\lambda=1}^{3}\{A_{\mathbf{0}}^{\lambda}\neq 0\}$, we employ the affine coordinates $z_{i}:=Z_{i}/Z_{0}$ for $i=1,\dots,5$ and $a_{\alpha}^{\lambda}:=A_{\alpha_{0}\alpha_{1}\alpha_{2}}^{\lambda}/A_{\mathbf{0}}^{\lambda}$ with $\alpha_{0}=d_{\lambda}-\alpha_{1}-\alpha_{2}$. To streamline notation, for each jet of order $j=1,2$ we define:
$\xi_{i}^{(j)}:=\begin{cases}z_{i}^{(j)}&\text{for }i=1,2,\\ (\mathrm{log}\,z_{2+\lambda})^{(j)}&\text{for }i=3,4,5\ (\lambda=i-2).\end{cases}$
Then local coordinates for $\overline{J_{2}}\big(\{Z_{0}\neq 0\}\times{\textstyle\prod}_{\lambda=1}^{3}\{A_{\mathbf{0}}^{\lambda}\neq 0\}\big)\subset J_{2}\big(\mathbb{P}^{5}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}+1},-\mathrm{log}\,\mathcal{D}\big)$ are given by the coordinate tuple $\left(z_{i},a_{\alpha}^{\lambda},\xi_{i_{1}}^{\prime},\xi_{i_{2}}^{\prime\prime}\right)$.
Following the strategy of Siu [Siu2004] and the realizations of Păun [mihaipaun2008] and Rousseau [Rousseau2009], we consider the following family of Siu’s slanted vector fields with pole order $7$:
Theorem 5.1.
Let $\Sigma_{0}$ be the zero locus of the $2\times 2$ Wronskian $\det(\xi_{i}^{(j)})_{1\leqslant i,j\leqslant 2}$:
$\Sigma_{0}=\left\{\left(z_{i},a_{\alpha}^{\lambda},\xi_{i_{1}}^{\prime},\xi_{i_{2}}^{\prime\prime}\right)\,\middle|\,\det\left(\xi_{i}^{(j)}\right)_{1\leqslant i,j\leqslant 2}=0\right\}\subset\overline{J_{2}^{v}}\left(\mathcal{Y}\cap\left(\{Z_{0}\neq 0\}\times{\textstyle\prod}_{\lambda=1}^{3}\{A_{\mathbf{0}}^{\lambda}\neq 0\}\right)\right),$
and $\Sigma$ be the closure of $\Sigma_{0}$ in $\overline{J_{2}^{v}}(\mathcal{Y}^{*})$. Then the twisted tangent bundle
(5.4) $T_{\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)}\otimes p^{*}\mathrm{pr}_{1}^{*}\mathcal{O}_{\mathbb{P}^{2}}(7)\otimes p^{*}\mathrm{pr}_{2}^{*}\mathcal{O}_{\prod_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}}(1,1,1),$
is generated by its global sections at every point of $\overline{J_{2}^{v}}(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}})\setminus\left(\pi_{[2]}\left(\Sigma\right)\cup p^{-1}\left(\mathcal{X}\right)\right)$, where $\pi_{[2]}$ is introduced in (5.3), $p$ is the natural projection
$p\colon\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)\longrightarrow\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}},$
and $\mathrm{pr}_{1}$ (resp. $\mathrm{pr}_{2}$) is the projection from $\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$ to $\mathbb{P}^{2}$ (resp. ${\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}$). ∎
5.2. Producing an Independent $2$-Jet via Siu’s Slanted Vector Fields of Pole Order 77
Proposition 5.2.
Given a holomorphic family $\{\omega_{1,a}\}_{a\in U}$ of negatively twisted invariant logarithmic $2$-jet differentials as in Proposition 4.4 with vanishing order $t\geqslant 8$, one can find some slanted vector field $V$ of pole order $7$ as a global section of (5.4), such that, the derivative of $\{\omega_{1,a}\}$ along $V$ gives a new family of holomorphic invariant logarithmic $2$-jet differentials $\omega_{2,a}$ with vanishing order $t-7\geqslant 1$. Furthermore, the common zero set $\{\omega_{1,a}=0\}\cap\{\omega_{2,a}=0\}$ in the second level $\left(\mathbb{P}^{2}_{a}\right)_{2}$ of the logarithmic Demailly–Semple tower of the log pair $\left(\mathbb{P}^{2},\mathcal{C}_{a}\right)$ has codimension at least $2$ modulo $\Gamma_{2}$ for a generic parameter $a\in U$.
This kind of proposition is due to Siu [Siu2004]. For the sake of completeness, we provide an argument here.
Proof.
We regard $\{\omega_{1,a}\}_{a\in U}$ as a holomorphic map
$P\colon\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{{\lambda}=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)\Big|_{\mathrm{pr}_{2}^{-1}\left(U\right)}\longrightarrow p^{*}\mathrm{pr}_{1}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-t)|_{\mathrm{pr}_{2}^{-1}\left(U\right)}.$
For any global section $V$ of (5.4), which is a meromorphic global vector field on $\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{\lambda=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)$ with values in $p^{*}\mathrm{pr}_{1}^{*}\mathcal{O}_{\mathbb{P}^{2}}(7)$, as shown in [mihaipaun2008, Rousseau2009], differentiating $P$ by $V$ yields a new holomorphic map:
$\mathrm{d}\,P(V)\colon\overline{J_{2}^{v}}\left(\mathbb{P}^{2}\times{\textstyle\prod}_{{\lambda}=1}^{3}\mathbb{P}^{N_{d_{\lambda}}}\right)\Big|_{\mathrm{pr}_{2}^{-1}\left(U\right)}\longrightarrow p^{*}\mathrm{pr}_{1}^{*}\mathcal{O}_{\mathbb{P}^{2}}(7-t)|_{\mathrm{pr}_{2}^{-1}\left(U\right)},$
which corresponds to a new holomorphic family of invariant logarithmic $2$-jet differentials with vanishing order $t-7\geqslant 1$.
By Theorem 5.1, the twisted tangent bundle (5.4) is globally generated outside $\pi_{[2]}\left(\Sigma\right)\cup p^{-1}\left(\mathcal{X}\right)$. Noting that $\{P|_{a}=\omega_{1,a}=0\}$ is irreducible and reduced modulo $\Gamma_{2}$, whence we can find some global section $V$ of (5.4) such that the intersection $\{P|_{a}=0\}\cap\{\mathrm{d}\,P(V)|_{a}=0\}$ has codimension 22 modulo $\Gamma_{2}$ for a generic parameter $a\in U$. ∎
6. Refinement of Demailly–El Goul’s Arguments
6.1. More Riemann-Roch Computations
In Proposition 4.4, we show the existence of a nonzero negatively twisted invariant logarithmic $2$-jet differential
$\omega_{1}\in H^{0}\big(X_{2},\mathcal{O}_{X_{2}}(m)\otimes\pi_{2,0}^{*}\mathcal{O}_{X}(-t)\big),$
for some positive integers $m,t\geqslant 1$ with $t/m\geqslant 1/5$, whose zero divisor on the second level of the Demailly-Semple tower $X_{2}$ is of the form $Z+b_{1}\Gamma_{2}$, where $Z\subset X_{2}$ is irreducible and reduced. Moreover, we have the linear equivalence
(6.1) $Z\sim b_{1}u_{1}+b_{2}u_{2}-t\pi_{2,0}^{*}h,\quad(b_{1},b_{2})\in\mathbb{Z}^{2},\quad b_{1}\geqslant 2b_{2}>0,\quad b_{1}+b_{2}=m,\quad t\in\mathbb{Z}_{+},$
where $u_{1}=\pi_{2,1}^{*}\mathcal{O}_{X_{1}}(1)$, $u_{2}=\mathcal{O}_{X_{2}}(1)$, and $h=\mathcal{O}_{\mathbb{P}^{2}}(1)$.
Hereafter, we use the standard notation:
$\mathcal{O}_{X_{2}}(b_{1},b_{2})\coloneqq\pi_{2,1}^{*}\mathcal{O}_{X_{1}}(b_{1})\otimes\mathcal{O}_{X_{2}}(b_{2}),\quad b_{1},b_{2}\in\mathbb{Z}.$
Using the same argument as in Demailly–El Goul [Demailly-Elgoul2000, Proposition 3.4] — an idea which originates in the work of Miyaoka [Miyaoka] and Lu–Yau [Lu-Yau] — one can show that $\mathcal{O}_{X_{2}}(2,1)|_{Z}$ is big, provided that $t<m$. Now, our goal is to determine an effective small rational number $\tau>0$ such that the $\mathbb{Q}$-line bundle $\big(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\big)|_{Z}$ remains big. This will be achieved through Riemann-Roch computations and the following proposition observed by Demailly [Demailly1997, p. 73].
Proposition 6.1.
For all integers $a_{2}>-2$ and $a_{1}>1$, the following higher direct images vanish:
$R^{q}(\pi_{2,0})_{*}\mathcal{O}_{X_{2}}(a_{1},a_{2})=0,\quad\forall q\geqslant 1.$
Since no readily available reference for this result exists in the literature, we provide a proof here.
Proof.
We first show that each fiber of $\pi_{2,0}:X_{2}\to X_{0}$ is isomorphic to the Hirzebruch surface (cf. [Beauville1996, p. 40])
$\mathbb{F}_{3}=\mathbb{P}\bigl(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(3)\bigr).$
Let $x\in X_{0}$ be an arbitrary point and denote by $F_{1}\cong\mathbb{P}^{1}$ the fiber of $\pi_{1,0}:X_{1}\to X_{0}$ over $x$. From the construction of the logarithmic Demailly-Semple tower (Subsection 2.2) we know that the fiber $F_{2}$ of $\pi_{2,0}$ over $x$ satisfies
$F_{2}\cong\mathbb{P}\bigl(V_{1}|_{F_{1}}\bigr),$
and the exact sequence (2.5) gives $V_{1}|_{F_{1}}\cong T_{X_{1}/X_{0}}|_{F_{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)$.
Using the relative Euler sequence (2.6) and taking determinants, we obtain
$T_{X_{1}/X_{0}}\cong\pi_{1,0}^{*}\,\overline{K}_{X}^{*}\otimes\mathcal{O}_{X_{1}}(2).$
Restricting to $F_{1}$ yields $T_{X_{1}/X_{0}}|_{F_{1}}\cong\mathcal{O}_{\mathbb{P}^{1}}(2)$. Hence
$V_{1}|_{F_{1}}\cong\mathcal{O}_{\mathbb{P}^{1}}(2)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1),$
and we conclude that $F_{2}\cong\mathbb{F}_{3}$.
On such a fiber we identify the restrictions of the relevant line bundles:
  • $\mathcal{O}_{X_{2}}(1)|_{F_{2}}$ is the tautological bundle, which we denote by $\mathcal{O}_{\mathbb{F}_{3}}(\hbar)$.
  • $\pi_{2,1}^{*}\mathcal{O}_{X_{1}}(1)|_{F_{2}}\cong\mathcal{O}_{\mathbb{F}_{3}}(f)$, where $f$ is a fiber of $F_{2}\rightarrow F_{1}$.
Consequently,
$\mathcal{O}_{X_{2}}(a_{1},a_{2})|_{F_{2}}\cong\mathcal{O}_{\mathbb{F}_{3}}(a_{2}\hbar+a_{1}f).$
Recall that the canonical bundle of $\mathbb{F}_{3}$ is
$K_{\mathbb{F}_{3}}\cong\mathcal{O}_{\mathbb{F}_{3}}(-2\hbar+f).$
Therefore,
$\mathcal{O}_{\mathbb{F}_{3}}(a_{2}\hbar+a_{1}f)\cong\mathcal{O}_{\mathbb{F}_{3}}\bigl((a_{2}+2)\hbar+(a_{1}-1)f\bigr)\otimes K_{\mathbb{F}_{3}}.$
Note that the line bundle $\mathcal{O}_{\mathbb{F}_{3}}\bigl((a_{2}+2)\hbar+(a_{1}-1)f\bigr)$ is ample precisely when
$a_{2}+2>0\quad\text{and}\quad a_{1}-1>0.$
Under these conditions, the Kodaira Vanishing Theorem gives
$H^{q}\bigl(\mathbb{F}_{3},\,\mathcal{O}_{\mathbb{F}_{3}}(a_{2}\hbar+a_{1}f)\bigr)=0,\qquad\forall\,q\geqslant 1.$
Hence
$H^{q}\bigl(F_{2},\,\mathcal{O}_{X_{2}}(a_{1},a_{2})|_{F_{2}}\bigr)=0,\qquad\forall\,x\in X_{0},\ F_{2}=\pi_{2,0}^{-1}(x),\ q\geqslant 1.$
Finally, Grauert’s Direct Image Theorem implies that the vanishing holds for the direct images on $X_{0}$:
$R^{q}(\pi_{2,0})_{*}\mathcal{O}_{X_{2}}(a_{1},a_{2})=0,\qquad\forall\,q\geqslant 1.$
Proposition 6.2.
If $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m}),\tau\in\mathbb{Q}$, where
$\tau_{1}\left(\frac{t}{m}\right)=\frac{3}{2}\left(4-\frac{t}{m}\right)-\frac{\sqrt{3}}{2}\sqrt{32-8\frac{t}{m}+3\left(\frac{t}{m}\right)^{2}},$
then the restricted $\mathbb{Q}$-line bundle $\big(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\big)|_{Z}$ is big on $Z$.
Proof.
We begin by recalling some key intersection formulas on the Demailly–Semple tower. By [Elgoul2003, p. 478], we have the identities:
$u_{1}^{2}+\bar{c}_{1}u_{1}+\bar{c}_{2}=0\quad\text{and}\quad u_{2}^{2}+\bar{c}_{1}(V_{1})u_{2}+\bar{c}_{2}(V_{1})=0,$
where $\bar{c}_{1}(V_{1})=\bar{c}_{1}+u_{1}$ (see (2.4)) and $\bar{c}_{2}(V_{1})=\bar{c}_{2}-u_{2}^{2}=2\bar{c}_{2}+\bar{c}_{1}u_{1}$.
On the second level $X_{2}$ of the Demailly-Semple tower, the following identities hold:
$u_{1}^{4}=0,\quad u_{1}^{3}u_{2}=\bar{c}_{1}^{2}-\bar{c}_{2},\quad u_{1}^{2}u_{2}^{2}=\bar{c}_{2},\quad u_{1}u_{2}^{3}=\bar{c}_{1}^{2}-3\bar{c}_{2},\quad u_{2}^{4}=5\bar{c}_{2}-\bar{c}_{1}^{2}.$
Furthermore, for any pullback $F\in\pi_{2,0}^{*}\operatorname{Pic}(\mathbb{P}^{2})$, we have:
$u_{1}^{3}\cdot F=0,\quad u_{1}^{2}u_{2}\cdot F=-\bar{c}_{1}\cdot F,\quad u_{1}u_{2}^{2}\cdot F=0,\quad u_{2}^{3}\cdot F=0.$
A crucial part of our approach relies on the following additional identities:
$u_{1}^{2}\cdot F_{1}\cdot F_{2}=0,\quad u_{1}u_{2}\cdot F_{1}\cdot F_{2}=F_{1}\cdot F_{2},\quad u_{2}^{2}\cdot F_{1}\cdot F_{2}=-F_{1}\cdot F_{2},\qquad\forall\,F_{1},F_{2}\in\pi_{2,0}^{*}\operatorname{Pic}(\mathbb{P}^{2}).$
Then by direct computation, we can obtain
(6.2) $\left(2u_{1}+u_{2}-\tau\pi_{2,0}^{*}h\right)^{3}\cdot Z=m\left(13\overline{c}_{1}^{2}-9\overline{c}_{2}\right)+12t(\overline{c}_{1}\cdot h)+12m(\overline{c}_{1}\cdot h)\tau+3m\left(h^{2}\right)\tau^{2}+9\left(h^{2}\right)t\tau.$
In our setting of three generic conics, we have $\overline{c}_{1}=-3h$ and $\overline{c}_{2}=9h^{2}$. Substituting these into (6.2), together with $h^{2}=1$, yields:
(6.3) $\left(2u_{1}+u_{2}-\tau\pi_{2,0}^{*}h\right)^{3}\cdot Z=3\left(m\tau^{2}-3(4m-t)\tau+12(m-t)\right).$
This quadratic polynomial has two real roots:
(6.4) $\tau_{1}\left(\frac{t}{m}\right)=\frac{3}{2}\left(4-\frac{t}{m}\right)-\frac{\sqrt{3}}{2}\sqrt{32-8\frac{t}{m}+3\left(\frac{t}{m}\right)^{2}}\,<\,\tau_{2}\left(\frac{t}{m}\right)=\frac{3}{2}\left(4-\frac{t}{m}\right)+\frac{\sqrt{3}}{2}\sqrt{32-8\frac{t}{m}+3\left(\frac{t}{m}\right)^{2}},$
because $32-8\frac{t}{m}+3(\frac{t}{m})^{2}>0$. Moreover, $\tau_{1}(\frac{t}{m})>0$ if and only if $\frac{t}{m}<1$.
Thus when $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m})$, we have $(2u_{1}+u_{2}-\tau\pi_{2,0}^{*}h)^{3}\cdot Z>0$. Riemann-Roch Theorem then gives, for sufficiently divisible integer $p\gg 1$, that:
$h^{0}\big(Z,(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}|_{Z}\big)+h^{2}\big(Z,(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}|_{Z}\big)$
$\geqslant$ $\chi\big(Z,(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}|_{Z}\big)=\frac{\big(2u_{1}+u_{2}-\tau\pi_{2,0}^{*}h\big)^{3}\cdot Z}{3!}p^{3}+O(p^{2})\gg 1.$
Now we show that
$h^{2}\big(Z,(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}|_{Z}\big)=0$
for all sufficiently divisible $p\gg 1$.
Consider the short exact sequence
[…]
and tensor it with $(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}$. The associated long exact cohomology sequence implies the dimension estimate:
$h^{2}\big(Z,(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}|_{Z}\big)$ $\leqslant h^{2}\big(X_{2},(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big)$
$\quad+h^{3}\big(X_{2},\mathcal{O}_{X_{2}}(-Z)\otimes(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big).$
First, we prove the vanishing
(6.5) $h^{2}\big(X_{2},(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big)=0.$
Consider the direct image sheaf $(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}$ on $\mathbb{P}^{2}$. By Proposition 6.1, the higher direct images vanish: $R^{q}(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1))^{p}=0$ for $q\geqslant 1$. By the projection formula, we obtain:
$R^{q}(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\cong R^{q}(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1))^{p}\otimes\mathcal{O}_{\mathbb{P}^{2}}(-p\tau)=0.$
Hence, Leray’s spectral sequence implies
$H^{2}\big(X_{2},(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big)\cong H^{2}\big(\mathbb{P}^{2},(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big).$
Now, by the isomorphism (cf. [Demailly-Elgoul2000, Lemma 3.3 (c)])
$(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1))^{p}\cong E_{2,3p}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}),$
we have
$H^{2}\big(\mathbb{P}^{2},(\pi_{2,0})_{*}(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big)\cong H^{2}\big(\mathbb{P}^{2},E_{2,3p}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-p\tau)\big).$
The same argument as in the proof of Proposition 4.2 shows that
$H^{2}\big(\mathbb{P}^{2},E_{2,3p}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-p\tau)\big)=0.$
Indeed, in the setting of Proposition 4.2, taking $m=3p$ and $\delta=\tau/9$, the condition $0\leqslant\delta<1/3$ is satisfied because $0<\frac{t}{m}<1$ and $0\leqslant\tau<\tau_{1}(\frac{t}{m})$ while the later is bounded from above by $6-2\sqrt{6}\approx 1.101$ by elementary calculation. This establishes (6.5).
Next, we prove the vanishing
(6.6) $H^{3}\big(X_{2},\mathcal{O}_{X_{2}}(-Z)\otimes(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big)=0.$
By (6.1), we have the isomorphism
$\mathcal{O}_{X_{2}}(-Z)\otimes\left(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\right)^{p}\cong\mathcal{O}_{X_{2}}(2p-b_{1},\,p-b_{2})\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(t-\tau p),$
with the inequality $2p-b_{1}\leqslant 2(p-b_{2})$.
As before, by Proposition 6.1, for all sufficiently divisible $p\gg 1$, the higher direct images vanish:
$R^{q}(\pi_{2,0})_{*}\left(\mathcal{O}_{X_{2}}(-Z)\otimes(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\right)=0\quad(q\geqslant 1).$
Thus, by Leray’s spectral sequence, we conclude that for all sufficiently divisible $p\gg 1$:
$H^{3}\big(X_{2},\mathcal{O}_{X_{2}}(-Z)\otimes(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau))^{p}\big)$
[…]
$=\$$0\qquad\text{[because\,\,\, $\dim\mathbb{P}^{2}=2<3$]}.$
This establishes (6.6), which completes the proof. ∎
6.2. Geometric Interpretation of Wronskians
Let $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$ be a normal crossing divisor of three smooth conics in $\mathbb{P}^{2}$, defined by homogeneous quadratic polynomials $A(\mathbf{Z})$, $B(\mathbf{Z})$, and $C(\mathbf{Z})$ in $\mathbb{C}[Z_{0},Z_{1},Z_{2}]$, where $\mathbf{Z}=(Z_{0},Z_{1},Z_{2})$.
On the affine chart $U_{0}\coloneqq\{Z_{0}\neq 0\}\subset\mathbb{P}^{2}$ with coordinates $(x_{1},x_{2})=(Z_{1}/Z_{0},Z_{2}/Z_{0})$, let $a(x_{1},x_{2})$, $b(x_{1},x_{2})$, and $c(x_{1},x_{2})$ be the dehomogenized polynomials of $A,B,C$ (see Subsection 7.1). Denote $a_{i}=\partial a/\partial x_{i}$ and $b_{j}=\partial b/\partial x_{j}$ for $i,j=1,2$.
Define the Zariski-open sets:
$U_{0,C}\coloneqq U_{0}\cap\{c(a_{1}b_{2}-b_{1}a_{2})\neq 0\},\,\,U_{0,B}\coloneqq U_{0}\cap\{b(a_{1}c_{2}-c_{1}a_{2})\neq 0\},\,\,U_{0,A}\coloneqq U_{0}\cap\{a(b_{1}c_{2}-c_{1}b_{2})\neq 0\}.$
For three generic smooth conics, we have
$\{c(a_{1}b_{2}-b_{1}a_{2})=0\}\cap\{b(a_{1}c_{2}-c_{1}a_{2})=0\}\cap\{a(b_{1}c_{2}-c_{1}b_{2})=0\}=\varnothing,$
so that $U_{0}=\cup_{Q\in\{A,B,C\}}U_{0,Q}$. Similarly, we define Zariski-open sets $U_{i,Q}$ for $i=1,2$ and $Q\in\{A,B,C\}$. These sets $U_{i,Q}$ form an open covering of $\mathbb{P}^{2}$.
On $U_{0,C}$, noting that $c$ is nowhere vanishing, the logarithmic $1$-forms $(\mathrm{log}\,a)^{\prime}$ and $(\mathrm{log}\,b)^{\prime}$ generate the logarithmic cotangent bundle $T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$. Let $f\colon(\mathbb{C},0)\ni t\mapsto f(t)\in U_{0,C}\subset\mathbb{P}^{2}$ be a germ of a holomorphic curve, with local representation $f=(f_{1},f_{2})$ in the coordinates $(x_{1},x_{2})$. Denote
(6.7) $a(f(t))\coloneqq a(f_{1}(t),f_{2}(t))\;\text{ and }\;b(f(t))\coloneqq b(f_{1}(t),f_{2}(t)).$
Consider the Wronskian operator
(6.8) $\begin{vmatrix}(\mathrm{log}\,a(f))^{\prime}&(\mathrm{log}\,b(f))^{\prime}\\ (\mathrm{log}\,a(f))^{\prime\prime}&(\mathrm{log}\,b(f))^{\prime\prime}\end{vmatrix}$.
Under a biholomorphic reparametrization $\varphi\colon(\mathbb{C},0)\rightarrow(\mathbb{C},0)$, the chain rule gives:
$(\mathrm{log}\,a(f\circ\varphi))^{\prime}$ $=\varphi^{\prime}\cdot(\mathrm{log}\,a(f))^{\prime}\circ\varphi,$
$(\mathrm{log}\,a(f\circ\varphi))^{\prime\prime}$ $=(\varphi^{\prime})^{2}\cdot(\mathrm{log}\,a(f))^{\prime\prime}\circ\varphi+\varphi^{\prime\prime}\cdot(\mathrm{log}\,a(f))^{\prime}\circ\varphi,$
with analogous expressions for $(\mathrm{log}\,b(f\circ\varphi))^{\prime}$ and $(\mathrm{log}\,b(f\circ\varphi))^{\prime\prime}$. A direct computation then shows:
$\begin{vmatrix}(\mathrm{log}\,a(f\circ\varphi))^{\prime}&(\mathrm{log}\,b(f\circ\varphi))^{\prime}\\ (\mathrm{log}\,a(f\circ\varphi))^{\prime\prime}&(\mathrm{log}\,b(f\circ\varphi))^{\prime\prime}\end{vmatrix}=(\varphi^{\prime})^{3}\cdot\begin{vmatrix}(\mathrm{log}\,a(f))^{\prime}&(\mathrm{log}\,b(f))^{\prime}\\ (\mathrm{log}\,a(f))^{\prime\prime}&(\mathrm{log}\,b(f))^{\prime\prime}\end{vmatrix}\circ\varphi.$
Hence, by Definition (2.2), the Wronskian operator (6.8) defines an invariant logarithmic $2$-jet differential on $U_{0,C}$ of weighted degree $3$.
Proposition 6.3.
Over the Zariski-open subset $\pi_{2,0}^{-1}(U_{0,C})$, the Wronskian operator (6.8) corresponds to a nowhere vanishing holomorphic section $w_{A,B,0}$ of the line bundle $\mathcal{O}_{X_{2}}(2,1)$.
While such properties are classical in the compact case (see, e.g., [Demailly1997]), to our knowledge, no convenient reference covers the logarithmic setting with three generic conics. For this reason, we include an explicit proof.
Proof.
Since the Wronskian operator (6.8) defines an invariant logarithmic $2$-jet differential of weighted degree $3$ on $U_{0,C}$, the direct image formula
$(\pi_{2,0})_{*}\mathcal{O}_{X_{2}}(m)\cong\mathcal{O}(E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C}))$
implies that it corresponds to a nonzero section $w_{A,B,0}$ of $\mathcal{O}_{X_{2}}(3)$ over $U_{0,C}$.
We now determine the vanishing order of $w_{A,B,0}$ along $\Gamma_{2}$ (see (2.7)). For any $(x_{1},x_{2})\in U_{0,C}$, we construct a germ $f\colon(\mathbb{C},0)\rightarrow(U_{0,C},(x_{1},x_{2}))$ of a logarithmic singular curve satisfying
$(f_{1}(0),f_{2}(0))=(x_{1},x_{2})\quad\text{and}\quad((\mathrm{log}\,a(f))^{\prime}(0),(\mathrm{log}\,b(f))^{\prime}(0))=(0,0),$
by recovering $(f_{1},f_{2})$ from the ODE system:
(6.9) $\begin{pmatrix}(\mathrm{log}\,a(f))^{\prime}(t)\\ (\mathrm{log}\,b(f))^{\prime}(t)\end{pmatrix}=\begin{pmatrix}a_{1}/a&a_{2}/a\\ b_{1}/b&b_{2}/b\end{pmatrix}\begin{pmatrix}f_{1}^{\prime}(t)\\ f_{2}^{\prime}(t)\end{pmatrix}=\begin{pmatrix}th_{a}(t)\\ th_{b}(t)\end{pmatrix}$,
where $h_{a}$ and $h_{b}$ are holomorphic functions to be chosen. Solving gives:
$\begin{pmatrix}f_{1}^{\prime}(t)\\ f_{2}^{\prime}(t)\end{pmatrix}=t\cdot\begin{vmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{vmatrix}^{-1}\begin{pmatrix}ab_{2}h_{a}(t)-ba_{2}h_{b}(t)\\ -ab_{1}h_{a}(t)+ba_{1}h_{b}(t)\end{pmatrix}$.
Since the middle determinant is nonvanishing on $U_{0,C}$, the holomorphic functions $(f_{1},f_{2})$ are determined by integration from the initial condition $(f_{1}(0),f_{2}(0))=(x_{1},x_{2})$.
Substituting (6.9) into the Wronskian (6.8) yields:
$\begin{vmatrix}(\mathrm{log}\,a(f))^{\prime}(t)&(\mathrm{log}\,b(f))^{\prime}(t)\\ (\mathrm{log}\,a(f))^{\prime\prime}(t)&(\mathrm{log}\,b(f))^{\prime\prime}(t)\end{vmatrix}=\begin{vmatrix}th_{a}(t)&th_{b}(t)\\ h_{a}(t)+th_{a}^{\prime}(t)&h_{b}(t)+th_{b}^{\prime}(t)\end{vmatrix}=t^{2}\begin{vmatrix}h_{a}(t)&h_{b}(t)\\ h_{a}^{\prime}(t)&h_{b}^{\prime}(t)\end{vmatrix}$.
By choosing $h_{a}$, $h_{b}$ such that the last determinant is nonvanishing, we see that $w_{A,B,0}$ vanishes to order exactly 22 along $\Gamma_{2}$. Therefore, by (2.8), we can regard $w_{A,B,0}$ as a nowhere vanishing section of
(6.10) $\mathcal{O}_{X_{2}}(3)\otimes\mathcal{O}(-2\Gamma_{2})\cong\mathcal{O}_{X_{2}}(2,1)$
on $\pi_{2,0}^{-1}(U_{0,C})$. ∎
6.3. A Refined Estimate
Here is a key step in proving our main theorem.
Theorem 6.4.
Let $g\colon\mathbb{C}\rightarrow\mathbb{P}^{2}$ be an entire curve whose lift $g_{[2]}\colon\mathbb{C}\rightarrow X_{2}$ satisfies $g_{[2]}(\mathbb{C})\subset Z$. Suppose there exists a section
(6.11) $\sigma\in H^{0}\Big(Z,\big(\mathcal{O}_{X_{2}}(2,1)^{\tilde{m}}\otimes\pi_{2,0}^{*}\mathcal{O}_{X}(-\tilde{t})\big)|_{Z}\Big)\cong H^{0}\Big(Z,\big(\mathcal{O}_{X_{2}}(3)\otimes\mathcal{O}(-2\Gamma_{2})\big)^{\tilde{m}}\otimes\pi_{2,0}^{*}\mathcal{O}_{X}(-\tilde{t})\big|_{Z}\Big),$
such that $\sigma(g_{[2]})\cdot(g_{[1]}^{\prime})^{3\tilde{m}}\not\equiv 0.$ Then the following Second Main Theorem type estimate holds:
$T_{g}(r)\leqslant\frac{2\tilde{m}}{\tilde{t}}\sum_{i=1}^{3}N_{g}^{(1)}(r,\mathcal{C}_{i})+o\big(T_{g}(r)\big)\quad\|.$
Our proof builds on the geometric interpretation of Wronskian operators and incorporates techniques from the proof of [Huynh-Vu-Xie-2017, Theorem 3.1]. This approach allows us to improve the constant on the right-hand side from $\frac{3\tilde{m}}{\tilde{t}}$ (cf. Theorem 1.2) to $\frac{2\tilde{m}}{\tilde{t}}$. This refinement constitutes a key contribution of the present work and underscores the advantage of using invariant logarithmic $2$-jet differentials over the Green-Griffiths ones.
Proof.
Consider the Wronskian operator (6.8) which corresponds to a nonvanishing section $w_{A,B,0}$ of $\mathcal{O}_{X_{2}}(2,1)$ on $\pi_{2,0}^{-1}(U_{0,C})$. Hence it induces a trivialization $\mathcal{O}_{X_{2}}\cdot w_{A,B,0}$ of $\mathcal{O}_{X_{2}}(2,1)$ on $\pi_{2,0}^{-1}(U_{0,C})$. Taking the $\tilde{m}-$power thus gives a trivialization $\mathcal{O}_{X_{2}}\cdot w_{A,B,0}^{\tilde{m}}$ of $\mathcal{O}_{X_{2}}(2,1)^{\tilde{m}}$ on $\pi_{2,0}^{-1}(U_{0,C})$.
Therefore, on $\pi_{2,0}^{-1}(U_{0,C})\cap Z$, we can write the section $\sigma$ as $\mathcal{R}\cdot w_{A,B,0}^{\tilde{m}}\cdot(\pi_{2,0}^{*}e)^{\tilde{t}}$ for some regular function $\mathcal{R}$ on $\pi_{2,0}^{-1}(U_{0,C})\cap Z$}, where $e$ is a local basis of $\mathcal{O}_{X}(-1)$ on $U_{0,C}$.
[…]
$\left\|g_{[2]}^{*}\sigma\cdot(g_{[1]}^{\prime})^{3\tilde{m}}\right\|_{g_{[2]}^{*}(\pi_{2,0}^{*}\mathcal{O}_{X}(-\tilde{t}))}$ $=\left|\left(g_{[2]}^{*}\mathcal{R}\right)\left(g_{[2]}^{*}w_{A,B,0}\right)^{\tilde{m}}\cdot\left(g_{[1]}^{\prime}\right)^{3\tilde{m}}\right|\left\|g_{[2]}^{*}\pi_{2,0}^{*}e\right\|_{g_{[2]}^{*}(\pi_{2,0}^{*}\mathcal{O}_{X}(-1))}^{\tilde{t}}$
[by $g=\pi_{2,0}\circ g_{[2]}$ and functoriality] $=\left|\left(g_{[2]}^{*}\mathcal{R}\right)\begin{vmatrix}\left(\mathrm{log}\,a(g)\right)^{\prime}&\left(\mathrm{log}\,b(g)\right)^{\prime}\\ \left(\mathrm{log}\,a(g)\right)^{\prime\prime}&\left(\mathrm{log}\,b(g)\right)^{\prime\prime}\end{vmatrix}^{\tilde{m}}\right|g^{*}\left\|e\right\|_{\mathcal{O}_{\mathbb{P}^{2}}(-1)}^{\tilde{t}}$.
Note that $g_{[2]}^{*}\mathcal{R}$ is a holomorphic function on $g^{-1}(U_{0,C}\cap\pi_{2,0}(Z))$.
Let $D_{\sigma,g_{[2]}}$ be the pole divisor of $g_{[2]}^{*}\sigma\cdot(g_{[1]}^{\prime})^{3\tilde{m}}$. Then by the Poincaré-Lelong formula, we obtain that
[…]
$\geqslant\tilde{t}g^{*}\omega_{\mathrm{FS}}-[D_{\sigma,g_{[2]}}],$
where functoriality guarantees that $g_{[2]}^{*}c_{1}\big(\pi_{2,0}^{*}\mathcal{O}_{X}(-1)\big)=-g^{*}\omega_{\mathrm{FS}}$.
Note that for $k=1,2$, the pole order of $\left(\mathrm{log}\,a(g)\right)^{(k)}$ (resp. $\left(\mathrm{log}\,b(g)\right)^{(k)}$) at $z\in g^{-1}(U_{0,C})\subset\mathbb{C}$ is at most $k\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{1},1\}$ (resp. $k\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{2},1\}$). Thus
$D_{\sigma,g_{[2]}}|_{g^{-1}(U_{0,C})}\leqslant\tilde{m}\Big(\sum_{z\in g^{-1}(U_{0,C})}2\big(\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{1},1\}+\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{2},1\}\big)\,z\Big).$
Note that $X_{2}$ can be covered by $\pi_{2,0}^{-1}(U_{i,Q})$ for $i=0,1,2$ and $Q\in\{A,B,C\}$. Applying similar argument on $U_{i,Q}$, we can obtain analogous pole order estimate on $g^{-1}(U_{i,Q})$. Combining these results, we get
$D_{\sigma,g_{[2]}}\leqslant 2\tilde{m}\sum_{z\in\mathbb{C}}\Big(\sum_{i=1}^{3}\min\{\operatorname{ord}_{z}g^{*}\mathcal{C}_{i},1\}\Big)z.$
Substituting this in the inequality (6.12) and applying Jensen’s formula, we obtain
$\tilde{t}\cdot T_{g}(r)\leqslant\int_{\partial\mathbb{D}_{r}}\mathrm{log}\,\|g_{[2]}^{*}\sigma\cdot(g_{[1]}^{\prime})^{3\tilde{m}}\|_{{}_{g_{[2]}^{*}(\pi_{2,0}^{*}\mathcal{O}_{X}(-\tilde{t}))}}\frac{\mathrm{d}\,\theta}{2\pi}+2\tilde{m}\,\sum_{i=1}^{3}N_{g}^{(1)}(r,\mathcal{C}_{i})+O(1).$
Therefore, to conclude the proposition, we need to show that
(6.13) $\int_{\partial\mathbb{D}_{r}}\mathrm{log}\,\|g_{[2]}^{*}\sigma\cdot(g_{[1]}^{\prime})^{3\tilde{m}}\|_{{}_{g_{[2]}^{*}(\pi_{2,0}^{*}\mathcal{O}_{X}(-\tilde{t}))}}\frac{\mathrm{d}\,\theta}{2\pi}=o\big(T_{g}(r)\big)\quad\parallel.$
The remaining argument follows the same line as the proof of (3.3) in [Huynh-Vu-Xie-2017, Theorem 3.1] and we omit it here. Hence we conclude the proof. ∎
6.4. An Effective Constant $c$ in the Second Main Theorem for Three Conics
Suppose there exists a nonzero negatively twisted invariant logarithmic $2$-jet differential
$\omega_{1}\in H^{0}\bigl(\mathbb{P}^{2},\,E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\bigr)$
satisfying $0<t<m$ and such that the divisor $Z$ of $\omega_{1}$ on the second level $X_{2}$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C})$ is irreducible and reduced modulo $\Gamma_{2}$. Then, by the argument in Section 6.1, for any rational number $\tau\in(0,\tau_{1}(t/m))$, we obtain a nontrivial global section
$\sigma\in H^{0}\left(Z,\,\left(\mathcal{O}_{X_{2}}(2,1)\otimes\pi_{2,0}^{*}\mathcal{O}_{\mathbb{P}^{2}}(-\tau)\right)^{\otimes\ell}\right)$
for sufficiently divisible $\ell\gg 1$ (depending on $\tau$).
Furthermore, if $f^{*}\omega_{1}\equiv 0$ but $f_{[2]}^{*}\sigma\not\equiv 0$, then Theorem 6.4 yields the estimate (1.5) with constant (see (6.4)) $c=\frac{2}{\tau}>\frac{2}{\tau_{1}(t/m)}.$ By the arbitrariness of $\tau$, the estimate (1.5) holds for any constant $c>2/\tau_{1}(\frac{t}{m})$.
Direct computation shows that $\inf_{0<x<1}\frac{2}{\tau_{1}(x)}=\frac{3+\sqrt{6}}{3}.$ Thus $\frac{2}{\tau_{1}(t/m)}<c$ is equivalent to $c>\frac{3+\sqrt{6}}{3}$ and $0<\frac{t}{m}<\frac{2(3c^{2}-6c+1)}{3c(2c-1)}.$
Meanwhile, in Case 1, to ensure that $\omega_{2}\in H^{0}\big(\mathbb{P}^{2},\,E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(7-t)\big)$ satisfies $m/(t-7)\leqslant c$, the positive integers $m$ and $t$ associated to the first family $\omega_{1}$ must satisfy $t\geqslant\frac{m}{c}+7.$
Taking $c=19$, we obtain the desired estimate (1.5) for positive integer pairs $(m,t)$ satisfying $0<t<\frac{1940}{2109}m$ or $t\geqslant\frac{m}{19}+7.$ Note that $\frac{1940}{2109}\times 8\approx 7.36$ and $\frac{1940}{2109}\times 7<7$. To exclude the exceptional cases of $(m,t)$, it suffices to prove the Key Vanishing Lemma I.
Taking $c=5$, we obtain the desired estimate (1.5) for positive integer pairs $(m,t)$ satisfying $0<t<\frac{92}{135}m$ or $t\geqslant\frac{m}{5}+7$. One may verify that for $m\geqslant 14$, every $t\geqslant 1$ satisfies one of these inequalities, since $\left\lceil\frac{92}{135}m\right\rceil\geqslant\left\lceil\frac{m}{5}+7\right\rceil$ for $m\geqslant 14$.
For $3\leqslant m\leqslant 14$, the following table provides the relevant data.
$m$ 3 4 5 6 7 8 9 10 11 12 13 14
$\frac{92}{135}m\approx$ 2.04 2.73 3.41 4.09 4.77 5.45 6.13 6.81 7.50 8.18 8.86 9.54
$\left\lfloor\frac{92}{135}m\right\rfloor$ 2 2 3 4 4 5 6 6 7 8 8 9
Thus, to exclude the exceptional cases of $(m,t)$, we only need to establish the Key Vanishing Lemma II.
7. Proof of the Key Vanishing Lemmas
7.1. Basic Calculations
Let $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$ be a divisor of three smooth conics in $\mathbb{P}^{2}$, defined by homogeneous quadratic polynomials:
$A(Z)$ $=A_{200}Z_{0}^{2}+A_{020}Z_{1}^{2}+A_{002}Z_{2}^{2}+A_{110}Z_{0}Z_{1}+A_{101}Z_{0}Z_{2}+A_{011}Z_{1}Z_{2},$
$B(Z)$ $=B_{200}Z_{0}^{2}+B_{020}Z_{1}^{2}+B_{002}Z_{2}^{2}+B_{110}Z_{0}Z_{1}+B_{101}Z_{0}Z_{2}+B_{011}Z_{1}Z_{2},$
$C(Z)$ $=C_{200}Z_{0}^{2}+C_{020}Z_{1}^{2}+C_{002}Z_{2}^{2}+C_{110}Z_{0}Z_{1}+C_{101}Z_{0}Z_{2}+C_{011}Z_{1}Z_{2},$
where $Z=(Z_{0},Z_{1},Z_{2})$ and all coefficients $A_{\bullet},B_{\bullet},C_{\bullet}\in\mathbb{C}$ are generic.
On the affine chart $U_{0}=\{Z_{0}\neq 0\}$ with coordinates $(x_{1},x_{2})=(Z_{1}/Z_{0},Z_{2}/Z_{0})$, the dehomogenized polynomials are:
$a(x_{1},x_{2})$ $=A_{200}+A_{020}x_{1}^{2}+A_{002}x_{2}^{2}+A_{110}x_{1}+A_{101}x_{2}+A_{011}x_{1}x_{2},$
$b(x_{1},x_{2})$ $=B_{200}+B_{020}x_{1}^{2}+B_{002}x_{2}^{2}+B_{110}x_{1}+B_{101}x_{2}+B_{011}x_{1}x_{2},$
$c(x_{1},x_{2})$ $=C_{200}+C_{020}x_{1}^{2}+C_{002}x_{2}^{2}+C_{110}x_{1}+C_{101}x_{2}+C_{011}x_{1}x_{2}.$
Their partial derivatives are given by:
$a_{1}$ $:=\frac{\partial a}{\partial x_{1}}=2A_{020}x_{1}+A_{110}+A_{011}x_{2},\quad a_{2}:=\frac{\partial a}{\partial x_{2}}=2A_{002}x_{2}+A_{101}+A_{011}x_{1},$
$b_{1}$ $:=\frac{\partial b}{\partial x_{1}}=2B_{020}x_{1}+B_{110}+B_{011}x_{2},\quad b_{2}:=\frac{\partial b}{\partial x_{2}}=2B_{002}x_{2}+B_{101}+B_{011}x_{1},$
$c_{1}$ $:=\frac{\partial c}{\partial x_{1}}=2C_{020}x_{1}+C_{110}+C_{011}x_{2},\quad c_{2}:=\frac{\partial c}{\partial x_{2}}=2C_{002}x_{2}+C_{101}+C_{011}x_{1}.$
The Jacobian determinant is the homogeneous cubic polynomial:
$J(Z):=\begin{vmatrix}\frac{\partial}{\partial Z_{0}}A&\frac{\partial}{\partial Z_{0}}B&\frac{\partial}{\partial Z_{0}}C\\ \frac{\partial}{\partial Z_{1}}A&\frac{\partial}{\partial Z_{1}}B&\frac{\partial}{\partial Z_{1}}C\\ \frac{\partial}{\partial Z_{2}}A&\frac{\partial}{\partial Z_{2}}B&\frac{\partial}{\partial Z_{2}}C\end{vmatrix}$.
Its vanishing locus $\{J=0\}$ is the ramification divisor of the endormorphism:
(7.1) $\sigma\colon\mathbb{P}^{2}\rightarrow\mathbb{P}^{2},\quad[Z_{0}:Z_{1}:Z_{2}]\mapsto[A:B:C].$
The utility of such Jacobian curves has been well-documented in the literature, cf. e.g. [Francois-Duval-2001, Corvaja-Zannier-2008, Guo-Sun-Wang-2021, Guo-Sun-Wang-2022, CHSX2025].
Under generic coefficient assumptions, the Jacobian curve $\{J=0\}$ satisfies the following transversality properties (cf., e.g., [Guo-Sun-Wang-2022, CHSX2025]):
  • (TP1)
    Transverse intersection with each conic $\mathcal{C}_{i}$ for $i=1,2,3$.
  • (TP2)
    Transverse intersection with each coordinate hyperplane $\{Z_{i}=0\}$ for $i=0,1,2$.
For the Key Vanishing Lemma I (but not for II), we will work with parameters satisfying these transversality conditions.
We begin with the following two results, which follow from direct computation on the affine charts $\{Z_{i}\neq 0\}\subset\mathbb{P}^{2}$, $i=0,1,2$.
Proposition 7.1.
Let $\mathcal{H}:=\{Z_{0}=0\}+\{Z_{1}=0\}+\{Z_{2}=0\}$ be a divisor consisting of the three coordinate hyperplanes in $\mathbb{P}^{2}$. Then the logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{H})$ is globally generated by the sections $(\mathrm{log}\,\frac{Z_{0}}{Z_{2}})^{\prime}$ and $(\mathrm{log}\,\frac{Z_{1}}{Z_{2}})^{\prime}$. ∎
Proposition 7.2.
The logarithmic $2$-jet differential bundle $E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{H})$ of weight $3$ is globally generated by the following sections:
(7.2) $\left(\left(\mathrm{log}\,\frac{Z_{0}}{Z_{2}}\right)^{\prime}\right)^{k}\left(\left(\mathrm{log}\,\frac{Z_{1}}{Z_{2}}\right)^{\prime}\right)^{m-3j-k}\begin{vmatrix}\left(\mathrm{log}\,\frac{Z_{0}}{Z_{2}}\right)^{\prime}&\left(\mathrm{log}\,\frac{Z_{1}}{Z_{2}}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{Z_{0}}{Z_{2}}\right)^{\prime\prime}&\left(\mathrm{log}\,\frac{Z_{1}}{Z_{2}}\right)^{\prime\prime}\end{vmatrix}^{j}\quad(0\leqslant j\leqslant\lfloor m/3\rfloor,\;0\leqslant k\leqslant m-3j).$
The endomorphism $\sigma$ defined in (7.1) is in fact a morphism $\sigma:(\mathbb{P}^{2},\mathcal{C})\to(\mathbb{P}^{2},\mathcal{H})$ between two log pairs. The functoriality immediately yields the relations between logarithmic differential $1$-forms
$\Big(\mathrm{log}\,\frac{A}{C}\Big)^{\prime}=\sigma^{*}\Big(\mathrm{log}\,\frac{Z_{0}}{Z_{2}}\Big)^{\prime}\quad\text{and}\quad\Big(\mathrm{log}\,\frac{B}{C}\Big)^{\prime}=\sigma^{*}\Big(\mathrm{log}\,\frac{Z_{1}}{Z_{2}}\Big)^{\prime}.$
Combining this with Proposition 7.1, we obtain:
Proposition 7.3.
On the Zariski-open set $\{J\neq 0\}\subset\mathbb{P}^{2}$, the logarithmic 1-forms $(\mathrm{log}\,\frac{A}{C})^{\prime}$ and $(\mathrm{log}\,\frac{B}{C})^{\prime}$ generate the sheaf $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$.
We complement this conceptual proof with explicit computations. While the functorial approach is elegant, the detailed coordinate computations reveal specific algebraic structures that will play an essential role in our later analysis.
Proof.
The sections $\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}$ and $\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}$ are clearly holomorphic on $\{J\neq 0\}$. We prove generation by working on affine charts.
On $U_{0}=\{Z_{0}\neq 0\}$ with coordinates $(x_{1},x_{2})$, consider the transformation matrix between differentials:
$\begin{pmatrix}\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\end{pmatrix}=M\begin{pmatrix}x_{1}^{\prime}\\ x_{2}^{\prime}\end{pmatrix},\quad M=\begin{pmatrix}\frac{\partial}{\partial x_{1}}\mathrm{log}\,\frac{a}{c}&\frac{\partial}{\partial x_{2}}\mathrm{log}\,\frac{a}{c}\\ \frac{\partial}{\partial x_{1}}\mathrm{log}\,\frac{b}{c}&\frac{\partial}{\partial x_{2}}\mathrm{log}\,\frac{b}{c}\end{pmatrix}$.
A direct computation gives the determinant of $M$:
(7.3) $\det M=\frac{D}{abc},\quad\text{where}\quad D=\begin{vmatrix}a&b&c\\ a_{1}&b_{1}&c_{1}\\ a_{2}&b_{2}&c_{2}\end{vmatrix}=\frac{J}{2Z_{0}^{3}}.$
Thus the matrix $M$ is invertible on the Zariski open set $U_{0}\cap\{J\neq 0\}\cap\{ABC\neq 0\}$, with its inverse given explicitly by:
$M^{-1}=\frac{1}{D}\begin{pmatrix}a(cb_{2}-bc_{2})&b(ac_{2}-ca_{2})\\ a(bc_{1}-cb_{1})&c(ba_{1}-ab_{1})\end{pmatrix}$.
Hence inverting the relation (7.3) yields the explicit coordinate transformation:
(7.4) $\begin{pmatrix}x_{1}^{\prime}\\ x_{2}^{\prime}\end{pmatrix}=\frac{1}{D}\begin{pmatrix}a(cb_{2}-bc_{2})&b(ac_{2}-ca_{2})\\ a(bc_{1}-cb_{1})&c(ba_{1}-ab_{1})\end{pmatrix}\begin{pmatrix}\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\end{pmatrix}$.
This establishes that the logarithmic differential forms $\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}$ and $\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}$ generate the sheaf of logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ over the Zariski-open set $\{J\neq 0\}\cap\{ABC\neq 0\}$, as the transformation matrix consists of regular functions on this domain.
Furthermore, the differential $(\mathrm{log}\,c)^{\prime}$ can be expressed in terms of our generators:
$(\mathrm{log}\,c)^{\prime}$ $=\frac{c^{\prime}}{c}=\begin{pmatrix}\frac{c_{1}}{c}&\frac{c_{2}}{c}\end{pmatrix}\begin{pmatrix}x_{1}^{\prime}\\ x_{2}^{\prime}\end{pmatrix}$
[use (7.4)] $=\frac{1}{D}\begin{pmatrix}a\begin{vmatrix}c_{1}&b_{1}\\ c_{2}&b_{2}\end{vmatrix}&b\begin{vmatrix}a_{1}&c_{1}\\ a_{2}&c_{2}\end{vmatrix}\end{pmatrix}\begin{pmatrix}\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\end{pmatrix}$.
Thus the sections $\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}$ and $\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}$ continue to generate the logarithmic cotangent bundle $T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ in a neighborhood of $\{J\neq 0\}\cap\{C=0\}$ inside $\{J\neq 0\}$, as the transformation matrix remains regular when approaching this locus. The analogous argument holds on charts $U_{1}=\{Z_{1}\neq 0\}$ and $U_{2}=\{Z_{2}\neq 0\}$, completing the proof. ∎
Using Proposition 7.2 and considering the morphism $\sigma:(\mathbb{P}^{2},\mathcal{C})\to(\mathbb{P}^{2},\mathcal{H})$, by functoriality, we obtain:
Proposition 7.4.
The bundle $E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ is freely generated over $\{J\neq 0\}$ by the logarithmic jet differentials:
(7.5) $\left(\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}\right)^{k}\left(\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}\right)^{m-3j-k}\begin{vmatrix}\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}&\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{A}{C}\right)^{\prime\prime}&\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime\prime}\end{vmatrix}^{j}\quad(0\leqslant j\leqslant\lfloor m/3\rfloor,\;0\leqslant k\leqslant m-3j).\qed$
Let $P\in\Gamma\big(\{J\neq 0\},E_{2,m}T^{*}{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}(-t)\big)$. To check and verify the global extendability of $P$, it suffices to demonstrate that $P$ admits a holomorphic extension across the Jacobian divisor $\{J=0\}$. Note that the intersection $\{J=0\}\cap\{Z_{0}ABC=0\}$ constitutes a Zariski-closed subset of codimension $2$ in $\mathbb{P}^{2}$. By Hartog’s extension theorem, it is thus sufficient to extend $P$ to $\{J=0\}\cap\{Z_{0}ABC\neq 0\}$. This extension problem necessitates an analysis of the transition formula for $E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}(-t)$ from $\{J\neq 0\}$ to $\{Z_{0}ABC\neq 0\}$.
Our carefully chosen basis (7.5) induces a highly homogeneous representation of the logarithmic jet differentials, which substantially mitigates the complications arising from inhomogeneous coordinate transitions on affine charts.333 We also attempted two other natural frame systems on $\{Z_{0}A\neq 0\}$ and $\{Z_{1}B\neq 0\}$, equipped with the obvious local frames, and applied the same strategy. Unfortunately, the computational complexity “explodes” beyond the capacity of current standard computers.
To convey the core ideas, we demonstrate the proof of Key Vanishing Lemma I in the “advanced case” $(m,t)=(3,1)$. The other cases, while verifiable by similar means, are omitted here because they follow systematically from the Key Vanishing Lemma II.
7.2. Exemplary Case $m=3,t=1$
For any global section $P\in H^{0}(\mathbb{P}^{2},E_{2,3}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}(-1))$, after suitable normalization, we may view $P$ as a global section of $E_{2,3}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ vanishing along a divisor $\{\zeta=0\}$ for some fixed $\zeta\in H^{0}(\mathbb{P}^{2},\mathcal{O}(1))$. For simplicity, we take $\zeta:=Z_{2}$. On $\{J\neq 0\}$, by Proposition 7.4, $P$ has the local decomposition:
$P=R\begin{vmatrix}\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}&\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{A}{C}\right)^{\prime\prime}&\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime\prime}\end{vmatrix}+\sum_{k=0}^{3}R^{k,3-k}\left(\left(\mathrm{log}\,\frac{A}{C}\right)^{\prime}\right)^{k}\left(\left(\mathrm{log}\,\frac{B}{C}\right)^{\prime}\right)^{3-k},$
where, by Serre’s gaga principle [MR82175] applied to the algebraic vector bundle $E_{2,3}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$, the coefficients $R,R^{k,3-k}$ are algebraic rational functions on $\{J\neq 0\}$; consequently they can be written in the form
(7.6) $\frac{K}{J^{\ell}}\quad\text{with}\quad\begin{cases}K\in\mathbb{C}[Z_{0},Z_{1},Z_{2}]_{3\ell}\text{ homogeneous},\\ \ell\geqslant 0\text{ integer},\\ \deg K=\deg J^{\ell}=3\ell.\end{cases}$
The core technical advances of our approach involve two principal components:
  • (i)
    Employing algebro-geometric techniques to establish a uniform upper bound of $\ell$, thereby reducing the infinite-dimensional problem to a computationally tractable finite-dimensional linear algebra framework;
  • (ii)
    Implementing a verification scheme (executed in Maple) that confirms the identical vanishing $P\equiv 0$ through explicit computation with certain carefully chosen coefficients $A_{\bullet}$, $B_{\bullet}$, and $C_{\bullet}$.
The semi-continuity argument then guarantees the extension of this vanishing property to all generic coefficients $A_{\bullet}$, $B_{\bullet}$, $C_{\bullet}$.
On the affine chart $U_{0}\cap\{J\neq 0\}=\{Z_{0}\neq 0\}\cap\{J\neq 0\}$ with coordinates $(x_{1},x_{2})=(Z_{1}/Z_{0},Z_{2}/Z_{0})$, the section $P$ has the local expression:
(7.7) $P=R_{0}(x_{1},x_{2})\begin{vmatrix}\left(\mathrm{log}\,\frac{a}{c}\right)^{\prime}&\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime}\\ \left(\mathrm{log}\,\frac{a}{c}\right)^{\prime\prime}&\left(\mathrm{log}\,\frac{b}{c}\right)^{\prime\prime}\end{vmatrix}+\sum_{k=0}^{3}R_{0}^{k,3-k}(x_{1},x_{2})\left(\Big(\mathrm{log}\,\frac{a}{c}\Big)^{\prime}\right)^{k}\left(\Big(\mathrm{log}\,\frac{b}{c}\Big)^{\prime}\right)^{3-k},$
where the coefficients $R_{0},R_{0}^{k,3-k}\in\mathbb{C}(x_{1},x_{2})$ are rational functions whose denominators consist of powers of the discriminant $D$ defined in (7.3), and which vanish to order $1$ along the divisor $\{x_{2}=0\}$.
When restricted to $U_{0}\cap\{ABC\neq 0\}$, the sheaf $E_{2,3}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ becomes free with basis:
(7.8) $\begin{vmatrix}x_{1}^{\prime}&x_{2}^{\prime}\\ x_{1}^{\prime\prime}&x_{2}^{\prime\prime}\end{vmatrix}\quad\text{and}\quad(x_{1}^{\prime})^{k}(x_{2}^{\prime})^{3-k}\quad(0\leqslant k\leqslant 3).$
Consequently, the section $P$ vanishing along $\{Z_{2}=0\}$ with order $1$ can be rewritten as:
(7.9) $P=\widetilde{R}_{0}(x_{1},x_{2})\begin{vmatrix}x_{1}^{\prime}&x_{2}^{\prime}\\ x_{1}^{\prime\prime}&x_{2}^{\prime\prime}\end{vmatrix}+\sum_{k=0}^{3}\widetilde{R}_{0}^{k,3-k}(x_{1},x_{2})(x_{1}^{\prime})^{k}(x_{2}^{\prime})^{3-k},$
where $\widetilde{R}_{0},\widetilde{R}_{0}^{k,3-k}$ are regular on $U_{0}\cap\{ABC\neq 0\}$ and vanish to order $1$ along the divisor $\{x_{2}=0\}$.
Comparing the expressions (7.7) and (7.9) on $U_{0}\cap\{ABC\neq 0\}\cap\{J\neq 0\}$ gives the coefficient transformation:
(7.10) $\begin{pmatrix}\widetilde{R}_{0}\\ \widetilde{R}_{0}^{30}\\ \widetilde{R}_{0}^{21}\\ \widetilde{R}_{0}^{12}\\ \widetilde{R}_{0}^{03}\end{pmatrix}=\Phi\begin{pmatrix}R_{0}\\ R_{0}^{30}\\ R_{0}^{21}\\ R_{0}^{12}\\ R_{0}^{03}\end{pmatrix}=\begin{pmatrix}(abc)^{-1}D&\mathbf{0}_{1\times 4}\\ \Delta&\Psi\end{pmatrix}\begin{pmatrix}R_{0}\\ R_{0}^{30}\\ R_{0}^{21}\\ R_{0}^{12}\\ R_{0}^{03}\end{pmatrix}$,
where $\Phi$ is the $5\times 5$ transition matrix between bases (7.5) and (7.8), with:
  • $\Delta=(abc)^{-2}\delta=(abc)^{-2}(\delta^{30},\delta^{21},\delta^{12},\delta^{03})^{\mathsf{T}}$, where $\delta^{30},\delta^{21},\delta^{12},\delta^{03}$ are some polynomials in $\mathbb{C}[A_{\bullet},B_{\bullet},C_{\bullet}][x_{1},x_{2}]$. To avoid diverting from our current focus, we omit their explicit expressions here.
  • $\Psi$ being the $4\times 4$ submatrix with factorization:
    (7.11) $\Psi=\psi\cdot\mathrm{diag}\left(-\frac{1}{a^{3}c^{3}},-\frac{1}{a^{2}bc^{3}},-\frac{1}{ab^{2}c^{3}},-\frac{1}{b^{3}c^{3}}\right),$
    where the last term means a diagonal matrix with the specified entries, and where $\psi$ is the matrix:
    $\begin{pmatrix}\alpha_{1}^{3}&\alpha_{1}^{2}\beta_{1}&\alpha_{1}\beta_{1}^{2}&\beta_{1}^{3}\\ 3\alpha_{1}^{2}\alpha_{2}&\alpha_{1}(2\alpha_{2}\beta_{1}+\alpha_{1}\beta_{2})&\beta_{1}(2\beta_{2}\alpha_{1}+\beta_{1}\alpha_{2})&3\beta_{1}^{2}\beta_{2}\\ 3\alpha_{1}\alpha_{2}^{2}&\alpha_{2}(2\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1})&\beta_{2}(2\beta_{1}\alpha_{2}+\beta_{2}\alpha_{1})&3\beta_{1}\beta_{2}^{2}\\ \alpha_{2}^{3}&\alpha_{2}^{2}\beta_{2}&\alpha_{2}\beta_{2}^{2}&\beta_{2}^{3}\end{pmatrix}$
    with $\alpha_{i}=ac_{i}-a_{i}c$ and $\beta_{i}=bc_{i}-b_{i}c$ for $i=1,2$.
From equation (7.10), we immediately obtain the relation:
(7.12) $\widetilde{R}_{0}=(abc)^{-1}DR_{0}.$
Under our standing assumption that the Jacobian curve $\{J=0\}$ intersects the three smooth conics $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ transversally, the discriminant $D$ is coprime to $a$, $b$, and $c$ in $\mathbb{C}[x_{1},x_{2}]$ (cf. [Guo-Sun-Wang-2022, CHSX2025]).
By the form of $R_{0}$ given in (7.6), we can write:
$R_{0}=\frac{t_{0}}{D^{\ell}}$
where $\ell\geqslant 0$, $t_{0}\in\mathbb{C}[x_{1},x_{2}]$ is coprime to $D$, and $\deg t_{0}\leqslant 3\ell$.
Key Observation 7.13.
The exponent satisfies $\ell\leqslant 1$.
Proof.
Assume, for contradiction, that $\ell\geqslant 2$. Through equation (7.12), we obtain the representation
$\widetilde{R}_{0}=\frac{t_{0}}{abcD^{\ell-1}}.$
The denominator reveals that $\widetilde{R}_{0}$ must have poles on the Zariski-open set $U_{0}\cap\{ABC\neq 0\}$ along $\{D=0\}$. This conclusion follows from:
  • Our geometric transversality assumption that $\{J=0\}$ meets each conic $\mathcal{C}_{i}$ transversally
  • The consequent coprimality between $D$ and abcabc
However, this contradicts the required regularity of $\widetilde{R}_{0}$. We therefore conclude that $\ell\leqslant 1$ must hold. ∎
Thus we can write
(7.14) $R_{0}=\frac{r_{0}}{D},\quad\text{with }r_{0}\in\mathbb{C}[x_{1},x_{2}]\text{ and }\deg r_{0}\leqslant 3.$
The remaining coefficients satisfy the transformation:
$\begin{pmatrix}\widetilde{R}_{0}^{30}\\ \widetilde{R}_{0}^{21}\\ \widetilde{R}_{0}^{12}\\ \widetilde{R}_{0}^{03}\end{pmatrix}=(abc)^{-2}R_{0}\begin{pmatrix}\delta^{30}\\ \delta^{21}\\ \delta^{12}\\ \delta^{03}\end{pmatrix}+\Psi\begin{pmatrix}R_{0}^{30}\\ R_{0}^{21}\\ R_{0}^{12}\\ R_{0}^{03}\end{pmatrix}$.
Substituting $R_{0}=r_{0}/D$ into the above equation yields:
(7.15) $\begin{pmatrix}R_{0}^{30}\\ R_{0}^{21}\\ R_{0}^{12}\\ R_{0}^{03}\end{pmatrix}=\frac{1}{(abc)^{2}D}\Psi^{-1}\left((abc)^{2}D\begin{pmatrix}\widetilde{R}_{0}^{30}\\ \widetilde{R}_{0}^{21}\\ \widetilde{R}_{0}^{12}\\ \widetilde{R}_{0}^{03}\end{pmatrix}-r_{0}\begin{pmatrix}\delta^{30}\\ \delta^{21}\\ \delta^{12}\\ \delta^{03}\end{pmatrix}\right)$.
To analyze the pole orders of the rational functions $R_{0}^{30},R_{0}^{21},R_{0}^{12},R_{0}^{03}$ along $\{D=0\}$, we compute $\Psi^{-1}$ explicitly using the factorization in (7.11). The key observation is (see (7.3)):
$\begin{vmatrix}\alpha_{1}&\beta_{1}\\ \alpha_{2}&\beta_{2}\end{vmatrix}=abc^{2}\det M=cD,$
which leads to:
$\Psi^{-1}=\frac{-1}{D^{3}}\operatorname{diag}(a^{3},a^{2}b,ab^{2},b^{3})\begin{pmatrix}\beta{{}_{2}}^{3}&-\beta_{1}\beta{{}_{2}}^{2}&\beta{{}_{1}}^{2}\beta_{2}&-\beta{{}_{1}}^{3}\\ -3\alpha_{2}\beta{{}_{2}}^{2}&\beta_{2}\left(2\beta_{1}\alpha_{2}+\beta_{2}\alpha_{1}\right)&-\beta_{1}\left(2\beta_{2}\alpha_{1}+\beta_{1}\alpha_{2}\right)&3\beta{{}_{1}}^{2}\beta_{2}\\ 3\alpha{{}_{2}}^{2}\beta_{2}&-\alpha_{2}\left(2\alpha_{1}\beta_{2}+\alpha_{2}\beta_{1}\right)&\alpha_{1}\left(2\alpha_{2}\beta_{1}+\alpha_{1}\beta_{2}\right)&-3\beta_{1}\beta{{}_{2}}^{2}\\ -\alpha{{}_{2}}^{3}&\alpha_{1}\alpha{{}_{2}}^{2}&-\alpha{{}_{1}}^{2}\alpha_{2}&\alpha{{}_{1}}^{3}\end{pmatrix}$
by direct computations.
The pole structure of the left-hand side in equation (7.15) is governed by three components of its right-hand side:
  • The prefactor $\frac{1}{(abc)^{2}D}$ contributes a simple pole along $\{D=0\}$.
  • The inverse matrix $\Psi^{-1}=-\frac{1}{D^{3}}[\cdots]$ introduces a triple pole along $\{D=0\}$.
  • The vector $\Psi^{-1}(\delta^{30},\delta^{21},\delta^{12},\delta^{03})^{\mathsf{T}}=\frac{1}{D^{2}}(\cdots)^{\mathsf{T}}$ produces a double pole along $\{D=0\}$ as a direct (Maple) computation shows. The essential step in our argument is to bound the pole orders along $\{D=0\}$ by explicit (though not necessarily sharp) effective exponents. This translates the problem into a concrete computational one. Since the required computations remain within the reach of contemporary machines, it becomes solvable. The existence of such an exponent bound would be foreshadowed and predicted by the well-known finiteness of cohomology $\dim H^{0}(\cdots)<\infty$.
  • The numerator terms $(abc)^{2}D\widetilde{R}_{0}^{k,3-k}$ and $r_{0}\delta^{k,3-k}$ remain regular on $\{D=0\}\cap\{abc\neq 0\}\cap U_{0}$.
Therefore, our transversality assumption — $\{J=0\}$ meets both the conics $\mathcal{C}_{i}$ and coordinate hyperplanes transversally — ensures that:
(7.16) $R_{0}^{k,3-k}=\frac{r_{0}^{k,3-k}}{D^{3}},\quad\text{where }r_{0}^{k,3-k}\in\mathbb{C}[x_{1},x_{2}].$
Furthermore, the degree bound $\deg r_{0}^{k,3-k}\leqslant 9$ is naturally constrained by the cubic nature of $D$.
Substituting the expressions from (7.14) and (7.16) into (7.10) yields the transformed coefficients:
(7.17) $\begin{pmatrix}\widetilde{R}_{0}\\ \widetilde{R}_{0}^{30}\\ \widetilde{R}_{0}^{21}\\ \widetilde{R}_{0}^{12}\\ \widetilde{R}_{0}^{03}\end{pmatrix}=\frac{-1}{(abc)^{3}D^{3}}\begin{pmatrix}\widetilde{r}_{0}\\ \widetilde{r}_{0}^{30}\\ \widetilde{r}_{0}^{21}\\ \widetilde{r}_{0}^{12}\\ \widetilde{r}_{0}^{03}\end{pmatrix}$,
where the numerator polynomials are determined by:
(7.18) $\begin{pmatrix}\widetilde{r}_{0}\\ \widetilde{r}_{0}^{30}\\ \widetilde{r}_{0}^{21}\\ \widetilde{r}_{0}^{12}\\ \widetilde{r}_{0}^{03}\end{pmatrix}=\begin{pmatrix}-abcD^{3}r_{0}\\ \psi\cdot\operatorname{diag}(b^{3},ab^{2},a^{2}b,a^{3})\cdot\mathbf{r}-abcD^{2}r_{0}\mathbf{\delta}\end{pmatrix}$,
with $\mathbf{r}=(r_{0}^{30},r_{0}^{21},r_{0}^{12},r_{0}^{03})^{\mathsf{T}}$ and $\mathbf{\delta}=(\delta^{30},\delta^{21},\delta^{12},\delta^{03})^{\mathsf{T}}$.
The vanishing of $R_{0}$ and $R_{0}^{k,3-k}$ along $\{x_{2}=0\}$ to order $1$, together with the expressions in (7.14) and (7.16), forces the polynomials $r_{0}$ and $r_{0}^{k,3-k}$ to adopt the specific forms:
(7.19) $r_{0}=x_{2}\sum_{\begin{subarray}{c}p+q\leqslant 2\end{subarray}}\lambda_{p,q}x_{1}^{p}x_{2}^{q},\quad r_{0}^{k,3-k}=x_{2}\sum_{\begin{subarray}{c}p+q\leqslant 8\end{subarray}}\lambda_{p,q}^{k,3-k}x_{1}^{p}x_{2}^{q}\qquad(\lambda_{p,q},\lambda_{p,q}^{k,3-k}\in\mathbb{C}).$
The regularity conditions for $\widetilde{R}_{0}$ and $\widetilde{R}_{0}^{k,3-k}$ on $U_{0}\cap\{ABC\neq 0\}$, combined with their first-order vanishing along $\{x_{2}=0\}$, impose the following divisibility requirement by (7.17): the numerators $\widetilde{r}_{0}$ and $\widetilde{r}_{0}^{k,3-k}$ must be divisible by $x_{2}D^{3}$ in $\mathbb{C}[x_{1},x_{2}]$.
From (7.18) and (7.19), it is clear that $\widetilde{r}_{0}=-(abc)^{2}x_{2}D^{3}\left(\sum_{\begin{subarray}{c}p+q\leqslant 2\end{subarray}}\lambda_{p,q}x_{1}^{p}x_{2}^{q}\right)$ satisfies the required divisibility condition. It therefore suffices to verify the divisibility condition for the remaining components $\widetilde{r}_{0}^{k,3-k}$ with $k=0,1,2,3$, which are given explicitly by:
(7.20) $\begin{pmatrix}\widetilde{r}_{0}^{30}\\ \widetilde{r}_{0}^{21}\\ \widetilde{r}_{0}^{12}\\ \widetilde{r}_{0}^{03}\end{pmatrix}=\psi\cdot\operatorname{diag}(b^{3},ab^{2},a^{2}b,a^{3})\cdot\mathbf{r}-abcD^{2}r_{0}\mathbf{\delta},$
where $\mathbf{r}=(r_{0}^{30},r_{0}^{21},r_{0}^{12},r_{0}^{03})^{\mathsf{T}}$ and $\mathbf{\delta}=(\delta^{30},\delta^{21},\delta^{12},\delta^{03})^{\mathsf{T}}$.
We consider the complex vector space $V_{3}$ spanned by the $\binom{2+2}{2}+4\times\binom{8+2}{2}=186$ coefficients $\lambda_{p,q},\lambda_{p,q}^{k,3-k}$ from (7.19). The target space $\widetilde{V}_{3}$ parameterizes the coefficients of four $x_{2}$-divisible polynomials in $\mathbb{C}[x_{1},x_{2}]$ with degree at most $24$, having dimension $\dim\widetilde{V}_{3}=4\times\binom{24-1+2}{2}=1200$.
The expression (7.20) naturally defines a linear transformation $\mathcal{F}_{3}:V_{3}\to\widetilde{V}_{3}$. The additional $D^{3}$-divisibility condition (see (7.17)) specifies a subspace $W_{3}\subset\widetilde{V}_{3}$ consisting of elements that factor as $\widetilde{r}_{0}^{k,3-k}=x_{2}D^{3}q_{0}^{k,3-k}$ where $\deg q_{0}^{k,3-k}\leqslant 14$, giving $\dim W_{3}=4\times\binom{14+2}{2}=480$.
By dimensional counting, we compare:
$\dim\mathcal{F}_{3}(V_{3})+\dim W_{3}\leqslant 186+480=666<1200=\dim\widetilde{V}_{3}.$
Hence it is “highly likely” that $\mathcal{F}_{3}(V_{3})\cap W_{3}=\{0\}$.
Indeed, this conclusion can be confirmed through explicit computer-assisted computations for the following specific three conics:
$A(Z)=2Z_{0}^{2}+Z_{1}^{2}+Z_{2}^{2}+Z_{0}Z_{1},\quad B(Z)=Z_{0}^{2}+Z_{1}^{2}+2Z_{2}^{2}+Z_{0}Z_{2},\quad C(Z)=Z_{0}^{2}+2Z_{1}^{2}+Z_{2}^{2}+Z_{1}Z_{2}.$
Thus by semi-continuity argument, we establish the the vanishing result:
$H^{0}\big(\mathbb{P}^{2},E_{2,3}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}(-1)\big)=0$
for generic three conics $\mathcal{C}$ in $\mathbb{P}^{2}$.
7.3. Proof of the Key Vanishing Lemma II
In what follows, we will abbreviate the first-order jet transfer formulas from standard jets $(x_{1}^{\prime},x_{2}^{\prime})$ to logarithmic jets $((\mathrm{log}\,\frac{a}{c})^{\prime},(\mathrm{log}\,\frac{b}{c})^{\prime})$ as:
$\begin{aligned} \Big(\mathrm{log}\,\frac{a}{c}\Big)^{\prime}&=\alpha\,x_{1}^{\prime}+\beta\,x_{2}^{\prime},\\ \Big(\mathrm{log}\,\frac{b}{c}\Big)^{\prime}&=\gamma\,x_{1}^{\prime}+\delta\,x_{2}^{\prime},\end{aligned}\qquad\qquad\begin{aligned} \alpha&:=\frac{a_{1}c-ac_{1}}{ac},&\beta&:=\frac{a_{2}c-ac_{2}}{ac},\\ \gamma&:=\frac{b_{1}c-bc_{1}}{bc},&\delta&:=\frac{b_{2}c-bc_{2}}{bc}.\end{aligned}$
Because logarithmic jet differentials are by definition assumed to exist and to be holomorphic outside the divisor $\mathcal{C}=\{ABC=0\}$, the presence of $a$, $b$, $c$ (within the affine chart $\mathbb{C}^{2}\subset\mathbb{P}^{2}$) in denominators does not produce any singularity.
The only possible singularity—which clarifies why Proposition 7.4 holds only on the complement $\mathbb{P}^{2}\setminus\{J=0\}$—comes from inverting the transfer formulas from logarithmic jets back to standard jets:
(7.21) $x_{1}^{\prime}$ $=\frac{\delta}{\alpha\delta-\beta\gamma}\,\Big(\mathrm{log}\,\frac{a}{c}\Big)^{\prime}-\frac{\beta}{\alpha\delta-\beta\gamma}\,\Big(\mathrm{log}\,\frac{b}{c}\Big)^{\prime},$
$x_{2}^{\prime}$ $=-\frac{\gamma}{\alpha\delta-\beta\gamma}\,\Big(\mathrm{log}\,\frac{a}{c}\Big)^{\prime}+\frac{\alpha}{\alpha\delta-\beta\gamma}\,\Big(\mathrm{log}\,\frac{b}{c}\Big)^{\prime},$
and this requires the nonvanishing of the $2\times 2$ determinant (see (7.3)):
(7.22) $\alpha\delta-\beta\gamma=\det\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}=\det M=\frac{D}{abc},\quad\text{where}\quad D=\begin{vmatrix}a&b&c\\ a_{1}&b_{1}&c_{1}\\ a_{2}&b_{2}&c_{2}\end{vmatrix}=\frac{J}{2Z_{0}^{3}}.$
By a direct computation in the affine chart $\{Z_{0}\neq 0\}$, this determinant happens to be a nonzero multiple of the Jacobian determinant $J$.
The transfer of the logarithmic Wronskian,
$\overline{W}^{\,\prime\prime\prime}\coloneqq\begin{vmatrix}\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{A}{C}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime\prime}\end{vmatrix}=\begin{vmatrix}\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}&\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{a}{c}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime\prime}\end{vmatrix}$,
to standard affine jet coordinates is more delicate, and is captured in the following Key Observation.
Key Observation 7.23.
When expressed in standard affine jet coordinates $(x_{1}^{\prime},x_{2}^{\prime},x_{1}^{\prime\prime},x_{2}^{\prime\prime})$, the logarithmic Wronskian takes the form:
(7.24) $\overline{W}^{\,\prime\prime\prime}$ $=(\alpha\gamma-\beta\delta)\,\begin{vmatrix}x_{1}^{\prime}&x_{2}^{\prime}\\ x_{1}^{\prime\prime}&x_{2}^{\prime\prime}\end{vmatrix}+\kappa\,{x_{1}^{\prime}}^{3}+\lambda\,{x_{1}^{\prime}}^{2}x_{2}^{\prime}+\mu\,x_{1}^{\prime}{x_{2}^{\prime}}^{2}+\nu\,{x_{2}^{\prime}}^{3},$
where the cubic terms in the first-order jets $x_{1}^{\prime}$, $x_{2}^{\prime}$ have coefficients $\kappa$, $\lambda$, $\mu$, $\nu$ given in terms of the second-order derivatives
$a_{11},\ a_{12},\ a_{22},\qquad b_{11},\ b_{12},\ b_{22},\qquad c_{11},\ c_{12},\ c_{22},$
and the coefficients $\alpha$, $\beta$, $\gamma$, $\delta$ as follows:
$\kappa$ $=-\gamma\frac{a_{11}}{a}+\alpha\frac{b_{11}}{b}-(\alpha-\gamma)\frac{c_{11}}{c}+\alpha^{2}\gamma-\alpha\gamma^{2},$
$\lambda$ $=-\delta\frac{a_{11}}{a}-2\gamma\frac{a_{12}}{a}+\beta\frac{b_{11}}{b}+2\alpha\frac{b_{12}}{b}-(\beta-\gamma)\frac{c_{11}}{c}-2(\alpha-\gamma)\frac{c_{12}}{c}$
$\qquad\qquad+\alpha^{2}\delta+2\alpha\beta\gamma-2\alpha\gamma\delta-\beta\gamma^{2},$
$\mu$ $=-2\delta\frac{a_{12}}{a}-\gamma\frac{a_{22}}{a}+2\beta\frac{b_{12}}{b}+\alpha\frac{b_{22}}{b}-2(\beta-\delta)\frac{c_{12}}{c}-(\alpha-\gamma)\frac{c_{22}}{c}$
$\qquad\qquad+2\alpha\beta\delta-\alpha\delta^{2}+\beta^{2}\gamma-2\beta\gamma\delta,$
$\nu$ $=-\delta\frac{a_{22}}{a}+\beta\frac{b_{22}}{b}-(\beta-\delta)\frac{c_{22}}{c}$
$\qquad\qquad+\beta^{2}\delta-\beta\delta^{2}.$
Proof.
This follows by direct computation. ∎
According to the statement of Key Vanishing Lemma II, we only need to consider $m\leqslant 13$. Thus, on the affine chart $(x_{1},x_{2})\in\mathbb{C}^{2}\subset\mathbb{P}^{2}$ away from the Jacobian zero-locus:
$\{D=0\}=\{J=0\}\cap\{Z_{0}\neq 0\},$
any invariant logarithmic $2$-jet differential of weighted degree $m\leqslant 13$ takes the general form:
(7.25) ${\sf Jet}_{ABC}$ $=\sum_{i+j=m}\,{R_{i,j}}\,\Big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\Big)^{j}$
$\quad+\sum_{i+j=m-3}\,{S_{i,j}}\,\Big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\Big)^{j}\,\overline{W}^{\,\prime\prime\prime}$
$\quad+\sum_{i+j=m-6}\,{U_{i,j}}\,\Big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\Big)^{j}\,\big(\overline{W}^{\,\prime\prime\prime}\big)^{2}$
$\quad+\sum_{i+j=m-9}\,{V_{i,j}}\,\Big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\Big)^{j}\,\big(\overline{W}^{\,\prime\prime\prime}\big)^{3}$
$\quad+\sum_{i+j=m-12}\,{W_{i,j}}\,\Big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\Big)^{j}\,\big(\overline{W}^{\,\prime\prime\prime}\big)^{4},$
which remains valid up to $m=14$, and where terms containing powers of $\overline{W}^{\prime\prime\prime}$ exceeding $m/3$ are understood to vanish identically. Here, the coefficients $R_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $S_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $U_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $V_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $W_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$ should be holomorphic (without singularity) outside $\{D=0\}$.
The gaga principle then ensures that each $R_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $S_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $U_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $V_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $W_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$ is a rational expression whose denominator is a certain integer power of $D$. Our preliminary objective is to determine these integers.
To this end, we consider the transfer formulas from logarithmic jets to standard jets. The first-order jet transfer formulas are already given in (7.21). Using (7.24), we have
$W^{\prime\prime\prime}\coloneqq\begin{vmatrix}x_{1}^{\prime}&x_{2}^{\prime}\\ x_{1}^{\prime\prime}&x_{2}^{\prime\prime}\end{vmatrix}=\frac{1}{\alpha\gamma-\beta\delta}\,\overline{W}^{\,\prime\prime\prime}-\frac{1}{\alpha\gamma-\beta\delta}\Bigl(\kappa\,{x_{1}^{\prime}}^{3}+\lambda\,{x_{1}^{\prime}}^{2}x_{2}^{\prime}+\mu\,x_{1}^{\prime}{x_{2}^{\prime}}^{2}+\nu\,{x_{2}^{\prime}}^{3}\Bigr),$
where the cubic term $\kappa\,{x_{1}^{\prime}}^{3}+\lambda\,{x_{1}^{\prime}}^{2}x_{2}^{\prime}+\mu\,x_{1}^{\prime}{x_{2}^{\prime}}^{2}+\nu\,{x_{2}^{\prime}}^{3}$ can be expanded using (7.21) as:
$\begin{pmatrix}{x_{1}^{\prime}}^{3}&{x_{1}^{\prime}}^{2}x_{2}^{\prime}&x_{1}^{\prime}{x_{2}^{\prime}}^{2}&{x_{2}^{\prime}}^{3}\end{pmatrix}\begin{pmatrix}\kappa\\ \lambda\\ \mu\\ \nu\end{pmatrix}$
$=$ $\begin{pmatrix}\big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\big)^{3}&\big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\big)^{2}\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}&\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\big)^{2}&\big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\big)^{3}\end{pmatrix}$
$\quad\cdot\frac{1}{(\alpha\gamma-\beta\delta)^{3}}\begin{pmatrix}\delta^{3}&-\gamma\delta^{2}&\gamma^{2}\delta&-\gamma^{3}\\ -3\beta\delta^{2}&\alpha\delta^{2}+2\beta\gamma\delta&-2\alpha\gamma\delta-\beta\gamma^{2}&3\alpha\gamma^{2}\\ 3\beta^{2}\delta&-2\alpha\beta\delta-\beta^{2}\gamma&\alpha^{2}\delta+2\alpha\beta\gamma&-3\alpha^{2}\gamma\\ -\beta^{3}&\alpha\beta^{2}&-\alpha^{2}\beta&\alpha^{3}\end{pmatrix}\begin{pmatrix}\kappa\\ \lambda\\ \mu\\ \nu\end{pmatrix}$.
Let us examine this more closely. Define
$\begin{pmatrix}\delta^{3}&-\gamma\delta^{2}&\gamma^{2}\delta&-\gamma^{3}\\ -3\beta\delta^{2}&\alpha\delta^{2}+2\beta\gamma\delta&-2\alpha\gamma\delta-\beta\gamma^{2}&3\alpha\gamma^{2}\\ 3\beta^{2}\delta&-2\alpha\beta\delta-\beta^{2}\gamma&\alpha^{2}\delta+2\alpha\beta\gamma&-3\alpha^{2}\gamma\\ -\beta^{3}&\alpha\beta^{2}&-\alpha^{2}\beta&\alpha^{3}\end{pmatrix}\begin{pmatrix}\kappa\\ \lambda\\ \mu\\ \nu\end{pmatrix}\eqqcolon\begin{pmatrix}\tau_{3,0}\\ \tau_{2,1}\\ \tau_{1,2}\\ \tau_{0,3}\end{pmatrix}$.
Miracle Factorizations 7.26.
The four expressions:
$\tau_{3,0}\,$= $\,\delta^{3}\,\kappa-\gamma\,\delta^{2}\,\lambda+\gamma^{2}\,\delta\,\mu-\gamma^{3}\,\nu$
$\,=$ $\,\big(\alpha\,\delta-\beta\,\gamma\big)\,\big(\cdots\big),$
$\tau_{2,1}\,$=  $\,-3\,\beta\,\delta^{2}\,\kappa+\big(\alpha\,\delta^{2}+2\,\beta\,\gamma\,\delta\big)\,\lambda+\big(-2\,\alpha\,\gamma\,\delta-\beta\,\gamma^{2}\big)\,\mu+3\,\alpha\,\gamma^{2}\,\nu$
$\,=$ $\,\big(\alpha\,\delta-\beta\,\gamma\big)\,\big(\cdots\big),$
$\tau_{1,2}\,$=  $3\,\beta^{2}\,\delta\,\kappa+(-2\,\alpha\,\beta\,\delta-\beta^{2}\,\gamma\big)\,\lambda+\big(\alpha^{2}\,\delta+2\,\alpha\,\beta\,\gamma\big)\,\mu-3\,\alpha^{2}\,\gamma\,\nu$
$\,=$ $\,\big(\alpha\,\delta-\beta\,\gamma\big)\,\big(\cdots\big),$
$\tau_{0,3}\,$= $\,-\,\beta^{3}\,\kappa+\alpha\,\beta^{2}\,\lambda-\alpha^{2}\,\beta\,\mu+\alpha^{3}\,\nu$
$\,=$ $\,\big(\alpha\,\delta-\beta\,\gamma\big)\,\big(\cdots\big),$
are all multiples of $\alpha\delta-\beta\gamma$.
Proof.
This follows by direct computation, which yields the four missing factors $(\cdots)$. ∎
Therefore:
$\begin{pmatrix}\tau_{3,0}\\ \tau_{2,1}\\ \tau_{1,2}\\ \tau_{0,3}\end{pmatrix}\eqqcolon(\alpha\delta-\beta\gamma)\begin{pmatrix}\theta_{3,0}\\ \theta_{2,1}\\ \theta_{1,2}\\ \theta_{0,3}\end{pmatrix}$,
and finally:
(7.27) $\begin{vmatrix}x_{1}^{\prime}&x_{2}^{\prime}\\ x_{1}^{\prime\prime}&x_{2}^{\prime\prime}\end{vmatrix}=\frac{1}{\alpha\gamma-\beta\delta}\begin{vmatrix}\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}&\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{a}{c}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime\prime}\end{vmatrix}-\frac{1}{(\alpha\gamma-\beta\delta)^{3}}\sum_{i+j=3}\theta_{i,j}\,\big(\big(\mathrm{log}\,\frac{a}{c}\big)^{\prime}\big)^{i}\big(\big(\mathrm{log}\,\frac{b}{c}\big)^{\prime}\big)^{j}.$
Recall that our main assumption is that, in the standard affine coordinates $(x_{1},x_{2})\in\mathbb{C}^{2}$, the logarithmic $2$-jet differential ${\sf Jet}_{ABC}$ is holomorphic over $\mathbb{C}^{2}\setminus\{abc=0\}$. Therefore, it can be expressed as
(7.28) ${\sf Jet}_{ABC}$ $=\sum_{i+j=m}\,\overline{R_{i,j}}\,{x_{1}^{\prime}}^{i}{x_{2}^{\prime}}^{j}+\sum_{i+j=m-3}\,\overline{S_{i,j}}\,{x_{1}^{\prime}}^{i}{x_{2}^{\prime}}^{j}\,W^{\prime\prime\prime}$
$\quad+\sum_{i+j=m-6}\,\overline{U_{i,j}}\,{x_{1}^{\prime}}^{i}{x_{2}^{\prime}}^{j}\,\big(W^{\prime\prime\prime}\big)^{2}+\sum_{i+j=m-9}\,\overline{V_{i,j}}\,{x_{1}^{\prime}}^{i}{x_{2}^{\prime}}^{j}\,\big(W^{\prime\prime\prime}\big)^{3}$
$\quad+\sum_{i+j=m-12}\,\overline{W_{i,j}}\,{x_{1}^{\prime}}^{i}{x_{2}^{\prime}}^{j}\,\big(W^{\prime\prime\prime}\big)^{4},$
where terms containing powers of ${W}^{\prime\prime\prime}$ exceeding $m/3$ are understood to vanish identically, and where all coefficients $\overline{R}_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $\overline{S}_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $\overline{U}_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $\overline{V}_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $\overline{W}_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$ are holomorphic on $\mathbb{C}^{2}\setminus\{abc=0\}$, and hence by the gaga principle are rational functions whose denominators may contain only $a$, $b$, $c$ as factors.
Now we examine the expression (7.25) for the jet differential ${\sf Jet}_{ABC}$. By substituting the transfer formulas (7.21) and (7.27) into the alternative expression (7.28) and comparing corresponding terms, we observe that the coefficients $\{R_{i,j}\}_{i+j=m}$, $\{S_{i,j}\}_{i+j=m-3}$, $\{U_{i,j}\}_{i+j=m-6}$, $\{V_{i,j}\}_{i+j=m-9}$, $\{W_{i,j}\}_{i+j=m-12}$ in (7.25) can be written as rational expressions whose denominators consist of factors $a$, $b$, $c$ (which are coprime to $D$) and $D$ (see (7.22)), but whose maximal possible pole orders with respect to $D$ are bounded by:
[…]
Thus, by an argument analogous to that of Key Observation 7.13, the invariant logarithmic $2$-jet differential ${\sf Jet}_{ABC}$ in (7.25) of weighted order $m\leqslant 13$ must, in homogeneous coordinates, take the following form:
(7.29) ${\sf Jet}_{ABC}$ $=\sum_{i+j=m}\,\frac{R_{i,j}(Z_{0},Z_{1},Z_{2})}{(J(Z_{0},Z_{1},Z_{2}))^{m}}\,\Big(\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\Big)^{j}\,\begin{vmatrix}\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{A}{C}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime\prime}\end{vmatrix}^{0}$
$\quad+\sum_{i+j=m-3}\,\frac{S_{i,j}(Z_{0},Z_{1},Z_{2})}{(J(Z_{0},Z_{1},Z_{2}))^{m-2}}\,\Big(\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\Big)^{j}\,\begin{vmatrix}\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{A}{C}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime\prime}\end{vmatrix}^{1}$
$\quad+\sum_{i+j=m-6}\,\frac{U_{i,j}(Z_{0},Z_{1},Z_{2})}{(J(Z_{0},Z_{1},Z_{2}))^{m-4}}\,\Big(\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\Big)^{j}\,\begin{vmatrix}\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{A}{C}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime\prime}\end{vmatrix}^{2}$
$\quad+\sum_{i+j=m-9}\,\frac{V_{i,j}(Z_{0},Z_{1},Z_{2})}{(J(Z_{0},Z_{1},Z_{2}))^{m-6}}\,\Big(\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\Big)^{j}\,\begin{vmatrix}\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{A}{C}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime\prime}\end{vmatrix}^{3}$
$\quad+\sum_{i+j=m-12}\,\frac{W_{i,j}(Z_{0},Z_{1},Z_{2})}{(J(Z_{0},Z_{1},Z_{2}))^{m-8}}\,\Big(\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}\Big)^{i}\,\Big(\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\Big)^{j}\,\begin{vmatrix}\big(\mathrm{log}\,\frac{A}{C}\big)^{\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime}\\ \big(\mathrm{log}\,\frac{A}{C}\big)^{\prime\prime}&\big(\mathrm{log}\,\frac{B}{C}\big)^{\prime\prime}\end{vmatrix}^{4},$
where terms containing powers of $\overline{W}^{\prime\prime\prime}$ exceeding $m/3$ are understood to vanish identically, and where the homogeneous polynomials
$R_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\,S_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\,U_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\,V_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\,W_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}\in\mathbb{C}[Z_{0},Z_{1},Z_{2}]$
are of constant degrees
$d_{R}=\deg R_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\quad d_{S}=\deg S_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\quad d_{U}=\deg U_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\quad d_{V}=\deg V_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},\quad d_{W}=\deg W_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}},$
satisfying
$d_{R}-3m=d_{S}-3(m-2)=d_{U}-3(m-4)=d_{V}-3(m-6)=d_{W}-3(m-8)=:t,$
so that this ${\sf Jet}_{ABC}$ is a meromorphic (rational) section of
$E_{2,m}T_{\mathbb{P}^{2}}^{\ast}\big(\mathrm{log}\,\mathcal{C}\big)\otimes\mathcal{O}_{\mathbb{P}^{2}}(t).$
For ${\sf Jet}_{ABC}$ to be a genuine holomorphic logarithmic jet differential, the coefficients of the polynomials $R_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $S_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $U_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $V_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$, $W_{{\scriptscriptstyle{\bullet}},{\scriptscriptstyle{\bullet}}}$ must satisfy a certain (large) linear system. This system encodes the requirement that, when passing to any of the three affine charts $\{Z_{0}\neq 0\}$, $\{Z_{1}\neq 0\}$, or $\{Z_{2}\neq 0\}$, and after simplifying all rational fractions, all powers of the Jacobian $J$ in the denominators must vanish completely.
To streamline explicit computer-assisted computations, we adopt a strategic simplification by choosing specific coefficients $A_{\bullet},B_{\bullet},C_{\bullet}$ such that the three conics $\mathcal{C}$ are of Fermat type; for example:
(7.30) $A(Z)=2Z_{0}^{2}+Z_{1}^{2}+Z_{2}^{2},\quad B(Z)=Z_{0}^{2}+2Z_{1}^{2}+Z_{2}^{2},\quad C(Z)=Z_{0}^{2}+Z_{1}^{2}+2Z_{2}^{2}.$
This choice ensures that the Jacobian $J=32Z_{0}Z_{1}Z_{2}$ takes an especially simple form, which significantly facilitates the verification of divisibility conditions in Maple.
A minor drawback is that the transversality condition (TP2) is no longer satisfied. However, this can be readily overcome using Hartogs’ extension theorem. Specifically, we carry out the analysis by passing from $\{J\neq 0\}$ to both affine charts $\{Z_{0}ABC\neq 0\}$ and $\{Z_{2}ABC\neq 0\}$, and verify the required divisibility in each chart separately.
More precisely, we work in the two affine charts:
$\{Z_{0}\neq 0\}\qquad\text{and}\qquad\{Z_{2}\neq 0\},$
with the coordinate substitutions:
$(Z_{0},Z_{1},Z_{2})\longmapsto(1,x_{1},x_{2})\qquad\text{and}\qquad(Z_{0},Z_{1},Z_{2})\longmapsto(y_{0},y_{1},1).$
In these coordinates, $J$ reads as $32x_{1}x_{2}$ or $32y_{0}y_{1}$. Thus, in each chart, all powers of $x_{1}x_{2}$ or $y_{0}y_{1}$ that appear in the denominators of the coefficients—obtained by substituting the transfer formulas into (7.29)—must be completely canceled by corresponding factors in the numerators. This cancellation requirement produces the system of linear equations that we need to solve.
In the program listed in the Appendix, we formulate the necessary divisibility conditions by (powers of) $x_{1}x_{2}$ and by (powers of) $y_{0}y_{1}$, and then ask the computer to solve the resulting linear system. To simplify the computation, we work modulo a prime $p=5$. The absence of nontrivial solutions in the reduced system implies the same for the original integer system, while making the verification significantly easier.
Using the Maple program provided in Appendix A, we can verify all instances of Key Vanishing Lemma II. The parameters $1\leqslant m\leqslant 13$ and $t\geqslant 1$ in the program can be adjusted manually to test different cases.444 Detailed explanations of these computations are available at: https://www.imo.universite-paris-saclay.fr/~joel.merker/d-m.pdf For the most challenging case $(m,t)=(13,9)$, the system comprised:
Equations: $\quad\mathbf{35,378},$
Variables: $\quad\mathbf{12,550},$
and was solved in Maple with the following resources:
Computation Time: $\quad\mathbf{22,117}\text{ seconds},$
Memory Usage: $\quad\mathbf{994.27}\text{ MB}.$
Remark 7.5.
Our computer-assisted computations for generic Fermat-type three conics $\mathcal{C}$ yield, as a byproduct, the expected vanishing for all $(m,t)=(1,1),(2,1),(3,1),(4,1),(5,2)$. In contrast, a nonvanishing result occurs unexpectedly for $(m,t)=(5,1)$:
$\dim\,E_{2,5}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-1)=9.$
The emergence of this nonvanishing case actually lends support to the reliability of our computations: had vanishing occurred for every tested pair $(m,t)$, the uniformity itself would appear suspicious. That a nonvanishing section surfaces already for such small parameters $(m,t)=(5,1)$ in the Fermat-type three-conic setting is, nevertheless, quite unexpected.
Hence, in the case where the pullback of one of these nine negatively twisted logarithmic $2$-jet differentials vanishes, we may apply Theorem 1.2 directly. Otherwise, we can proceed via Păun’s Proportionality Lemma [Demailly-Elgoul2000] combined with McQuillan’s Theorem 1.1. In either situation, one obtains a Second Main Theorem for such Fermat-type three conics $\mathcal{C}$ with constant $5$. It should be noted that this result does not follow directly from our main theorem, as the “generic” assumption therein may exclude the specific Fermat-type configuration considered here. Moreover, our explicit computations enable us to identify concrete examples of such Fermat-type three conics $\mathcal{C}$ satisfying the Second Main Theorem with constant $5$, such as the selection (7.30). Complete details will be presented in a forthcoming paper.
Acknowledgements
We are grateful to Julien Duval for introducing us a decade ago to the problem of establishing a Second Main Theorem for three conics in $\mathbb{P}^{2}$. We also thank Junjiro Noguchi for sharing his private notes, and Duc Viet Vu for offering valuable insights from a pluripotential perspective.
S.-Y. Xie would like to thank Daniel Barlet, Frédéric Campana, Yunling Chen, Tao Cui, Hanlong Fang, Jie Liu, Mihai Păun, Erwan Rousseau, Yum-Tong Siu, Ruiran Sun, and Kang Zuo for inspiring discussions. S.-Y. Xie and L. Hou also wish to thank Tianyuan Mathematics Research Center for its excellent environment.
We also thank Yunling Chen, Alexandre Eremenko, and Yi C. Huang for their suggestions on an earlier version. We thank Steven Lu and Katsutoshi Yamanoi for helpful comments.
The results of this article were first presented by S.-Y. Xie at The Seminar on Several Complex Variables and Complex Geometry (held from July 27 to August 2, 2025), which was organized by Xiangyu Zhou, Yum-Tong Siu, and John Erik Fornæss.
Funding
S.-Y. Xie acknowledges partial support from the National Key R&D Program of China under Grants No. 2023YFA1010500 and No. 2021YFA1003100, and from the National Natural Science Foundation of China under Grants No. 12288201 and No. 12471081, as well as support from the Xiaomi Young Talents Program. D.T. Huynh is supported by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under the grant number 101.04-2025.19.
Appendix A Maple Code for the Key Vanishing Lemmas
restart: with(LinearAlgebra): with(Groebner):
m := 13;
twist := - m + 4;
p := 5;
dR := 3*(m-0) + twist:
dS := 3*(m-2) + twist:
dU := 3*(m-4) + twist:
dV := 3*(m-6) + twist:
dW := 3*(m-8) + twist:
AA := 2*T^2 + 1*x^2 + 1*y^2;
$B$ := 1*T^2 + 2*x^2 + 1*y^2;
$C$ := 1*T^2 + 1*x^2 + 2*y^2;
AT := diff(AA,T): BT := diff($B$,T): CT := diff($C$,T):
AX := diff(AA,X): BX := diff($B$,X): CX := diff($C$,X):
AY := diff(AA,Y): BY := diff($B$,Y): CY := diff($C$,Y):
ATT := diff(AA,T,T): ATX := diff(AA,T,X): ATY := diff(AA,T,Y):
AXX := diff(AA,X,X): AXY := diff(AA,X,Y): AYY := diff(AA,Y,Y):
BTT := diff($B$,T,T): BTX := diff($B$,T,X): BTY := diff($B$,T,Y):
BXX := diff($B$,X,X): BXY := diff($B$,X,Y): BYY := diff($B$,Y,Y):
CTT := diff($C$,T,T): CTX := diff($C$,T,X): CTY := diff($C$,T,Y):
CXX := diff($C$,X,X): CXY := diff($C$,X,Y): CYY := diff($C$,Y,Y):
MatDD := Matrix([[AT,BT,CT],
[AX,BX,CX],
[AY,BY,CY]]):
$D$ := factor(Determinant(MatDD));
Ap := AT*T1 + AX*X1 + AY*Y1:
Bp := BT*T1 + BX*X1 + BY*Y1:
Cp := CT*T1 + CX*X1 + CY*Y1:
App := + AT*T2 + AX*X2 + AY*Y2
+ ATT*T1*T1 + 2*ATX*T1*X1 + 2*ATY*T1*Y1 + AXX*X1*X1 + 2*AXY*X1*Y1 + AYY*Y1*Y1:
Bpp := + BT*T2 + BX*X2 + BY*Y2
+ BTT*T1*T1 + 2*BTX*T1*X1 + 2*BTY*T1*Y1 + BXX*X1*X1 + 2*BXY*X1*Y1 + BYY*Y1*Y1:
Cpp := + CT*T2 + CX*X2 + CY*Y2
+ CTT*T1*T1 + 2*CTX*T1*X1 + 2*CTY*T1*Y1 + CXX*X1*X1 + 2*CXY*X1*Y1 + CYY*Y1*Y1:
logACp := Ap/AA - Cp/$C$:
logBCp := Bp/$B$ - Cp/$C$:
logACpp := App/AA - Ap^2/AAA^2 - (Cpp/$C$ - Cp^2/CCC^2):
logBCpp := Bpp/$B$ - Bp^2/BB^2 - (Cpp/$C$ - Cp^2/CCC^2):
Carte_xy := {T = 1, X = x, Y = y,
T1 = 0, X1 = x1, Y1 = y1,
T2 = 0, X2 = x2, Y2 = y2}:
xyAA := subs(Carte_xy, AA):
xyBB := subs(Carte_xy, $B$):
xyCC := subs(Carte_xy, $C$):
xyDD := subs(Carte_xy, $D$):
xylogACp := subs(Carte_xy, logACp):
xylogBCp := subs(Carte_xy, logBCp):
xylogACpp := subs(Carte_xy, logACpp):
xylogBCpp := subs(Carte_xy, logBCpp):
Carte_tx := {T = t, X = x, Y = 1,
T1 = t1, X1 = x1, Y1 = 0,
T2 = t2, X2 = x2, Y2 = 0}:
txAA := subs(Carte_tx, AA):
txBB := subs(Carte_tx, $B$):
txCC := subs(Carte_tx, $C$):
txDD := subs(Carte_tx, $D$):
txlogACp := subs(Carte_tx, logACp):
txlogBCp := subs(Carte_tx, logBCp):
txlogACpp := subs(Carte_tx, logACpp):
txlogBCpp := subs(Carte_tx, logBCpp):
Jet_ABC := + (1/(T^X^Y)^(m-0))
* add(R[m-0-j,j] * logACp^(m-0-j) * logBCp^j *
(logACp*logBCpp-logACpp*logBCp)^0, j=0..m-0)
+ (1/(T^X^Y)^(m-2))
* add(S[m-3-j,j] * logACp^(m-3-j) * logBCp^j *
(logACp*logBCpp-logACpp*logBCp)^1, j=0..m-3)
+ (1/(T^X^Y)^(m-4))
* add(U[m-6-j,j] * logACp^(m-6-j) * logBCp^j *
(logACp*logBCpp-logACpp*logBCp)^2, j=0..m-6)
+ (1/(T^X^Y)^(m-6))
* add(V[m-9-j,j] * logACp^(m-9-j) * logBCp^j *
(logACp*logBCpp-logACpp*logBCp)^3, j=0..m-9)
+ (1/(T^X^Y)^(m-8))
* add(W[m-12-j,j] * logACp^(m-12-j) * logBCp^j *
(logACp*logBCpp-logACpp*logBCp)^4, j=0..m-12):
xyJet_ABC := factor(+ (1/(x*y)^(m-0))
* add(Rxy[m-0-j,j] * xylogACp^(m-0-j) * xylogBCp^j *
(xylogACp*xylogBCpp-xylogACpp*xylogBCp)^0, j=0..m-0)
+ (1/(x*y)^(m-2))
* add(Sxy[m-3-j,j] * xylogACp^(m-3-j) * xylogBCp^j *
(xylogACp*xylogBCpp-xylogACpp*xylogBCp)^1, j=0..m-3)
+ (1/(x*y)^(m-4))
* add(Uxy[m-6-j,j] * xylogACp^(m-6-j) * xylogBCp^j *
(xylogACp*xylogBCpp-xylogACpp*xylogBCp)^2, j=0..m-6)
+ (1/(x*y)^(m-6))
* add(Vxy[m-9-j,j] * xylogACp^(m-9-j) * xylogBCp^j *
(xylogACp*xylogBCpp-xylogACpp*xylogBCp)^3, j=0..m-9)
+ (1/(x*y)^(m-8))
* add(Wxy[m-12-j,j] * xylogACp^(m-12-j) * xylogBCp^j *
(xylogACp*xylogBCpp-xylogACpp*xylogBCp)^4, j=0..m-12)):
txJet_ABC := + (1/(t*x)^(m-0))
* add(Rtx[m-0-j,j] * txlogACp^(m-0-j) * txlogBCp^j *
(txlogACp*txlogBCpp-txlogACpp*txlogBCp)^0, j=0..m-0)
+ (1/(t*x)^(m-2))
* add(Stx[m-3-j,j] * txlogACp^(m-3-j) * txlogBCp^j *
(txlogACp*txlogBCpp-txlogACpp*txlogBCp)^1, j=0..m-3)
+ (1/(t*x)^(m-4))
* add(Utx[m-6-j,j] * txlogACp^(m-6-j) * txlogBCp^j *
(txlogACp*txlogBCpp-txlogACpp*txlogBCp)^2, j=0..m-6)
+ (1/(t*x)^(m-6))
* add(Vtx[m-9-j,j] * txlogACp^(m-9-j) * txlogBCp^j *
(txlogACp*txlogBCpp-txlogACpp*txlogBCp)^3, j=0..m-9)
+ (1/(t*x)^(m-8))
* add(Wtx[m-12-j,j] * txlogACp^(m-12-j) * txlogBCp^j *
(txlogACp*txlogBCpp-txlogACpp*txlogBCp)^4, j=0..m-12):
NFxy := proc(P,j,i)
expand(mtaylor(P,x,j) + mtaylor(P,y,i) - mtaylor(mtaylor(P,y,i), x,j)):
end proc:
NFtx := proc(P,j,i)
expand(mtaylor(P,t,j) + mtaylor(P,x,i) - mtaylor(mtaylor(P,x,i), t,j)):
end proc:
xyJet := mtaylor(
factor((x*y*xyAAA*xyBB*yyyCC)^m * (subs(y2=(xyW+x2*y1)/x1, xyJet_ABC)))
, [x1,y1,xyW], 100):
txJet := mtaylor(
factor((t*x*txAAA*txBB*txCC)^m * (subs(x2=(txW+t2*x1)/t1, txJet_ABC)))
, [t1,x1,txW], 100):
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
xyEEEq[m-j-3*k,j,k] := factor(coeftayl(xyJet, [x1,y1,xyW]={0,0,0}, [m-j-3*k,j,k])):
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
txEEEq[m-j-3*k,j,k] := factor(coeftayl(txJet, [t1,x1,txW]={0,0,0}, [m-j-3*k,j,k])):
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
xyEEEq[m-j-3*k,j,k] := mtaylor(
factor(
(1/(x^((m-j-3*k)*y^)*(x*y)^*(3*k))) *
(1/((xyAA*xyBB*xyCC)^*(2*k))) *
xyEEEq[m-j-3*k,j,k])
, [x,y], 1000):
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
txEEEq[m-j-3*k,j,k] := mtaylor(
factor(
(1/(t^m-j-3*k)*x^j*(t*x)^(3*k))) * txEEEq[m-j-3*k,j,k])
, [t,x], 1000):
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
xyEEq[m-j-3*k,j,k] := NFxy(xyEEEq[m-j-3*k,j,k], j,m-j-3*k) mod p:
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
txEEq[m-j-3*k,j,k] := NFtx(txEEEq[m-j-3*k,j,k], j,m-j-3*k) mod p:
od:
od:
for j from 0 to m-0 do
Rxy[m-0-j,j] := add(add(cR[m-0-j,j,dR-q-r,q,r]*1^dR-q-r)*x^q*y^r, q=0..dR-r), r=0..dR):
od:
for j from 0 to m-3 do
Sxy[m-3-j,j] := add(add(cS[m-3-j,j,dS-q-r,q,r]*1^dS-q-r)*x^q*y^r, q=0..dS-r), r=0..dS):
od:
for j from 0 to m-6 do
Uxy[m-6-j,j] := add(add(cU[m-6-j,j,dU-q-r,q,r]*1^dU-q-r)*x^q*y^r, q=0..dU-r), r=0..dU):
od:
for j from 0 to m-9 do
Vxy[m-9-j,j] := add(add(cV[m-9-j,j,dV-q-r,q,r]*t^((dV-q-r)*x^q*y^r, q=0..dV-r), r=0..dV):
od:
for j from 0 to m-12 do
Wxy[m-12-j,j] := add(add(cW[m-12-j,j,dW-q-r,q,r]*t^((dW-q-r)*x^q*y^r, q=0..dW-r), r=0..dW):
od:
for j from 0 to m-0 do
Rtx[m-0-j,j] := add(add(cR[m-0-j,j,dR-q-r,q,r]*t^((dR-q-r)*x^q*1^r, q=0..dR-r), r=0..dR):
od:
for j from 0 to m-3 do
Stx[m-3-j,j] := add(add(cS[m-3-j,j,dS-q-r,q,r]*t^((dS-q-r)*x^q*1^r, q=0..dS-r), r=0..dS):
od:
for j from 0 to m-6 do
Utx[m-6-j,j] := add(add(cU[m-6-j,j,dU-q-r,q,r]*t^((dU-q-r)*x^q*1^r, q=0..dU-r), r=0..dU):
od:
for j from 0 to m-9 do
Vtx[m-9-j,j] := add(add(cV[m-9-j,j,dV-q-r,q,r]*t^((dV-q-r)*x^q*1^r, q=0..dV-r), r=0..dV):
od:
for j from 0 to m-12 do
Wtx[m-12-j,j] := add(add(cW[m-12-j,j,dW-q-r,q,r]*t^((dW-q-r)*x^q*1^r, q=0..dW-r), r=0..dW):
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
xyEq[m-j-3*k,j,k] := NFxy(xyEEq[m-j-3*k,j,k], j,m-j-3*k):
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
txEq[m-j-3*k,j,k] := NFtx(txEEq[m-j-3*k,j,k], j,m-j-3*k):
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
xyEquations[m-j-3*k,j,k] := {coeffs(xyEq[m-j-3*k,j,k], [x,y])}:
od:
od:
printlevel := 2:
for k from 0 to m/3 do
for j from 0 to m-3*k do
txEquations[m-j-3*k,j,k] := {coeffs(txEq[m-j-3*k,j,k], [t,x])}:
od:
od:
xyEquations_Add := {}
for k from 0 to m/3 do
for j from 0 to m-3*k do
xyEquations_Add := xyEquations_Add union xyEquations[m-j-3*k,j,k]:
od:
od:
Equations_xy := xyEquations_Add:
txEquations_Add := {};
for k from 0 to m/3 do
for j from 0 to m-3*k do
txEquations_Add := txEquations_Add union txEquations[m-j-3*k,j,k]:
od:
od:
Equations_tx := txEquations_Add:
Equations := Equations_xy union Equations_tx mod p:
Nombre_Equations := nops(Equations);
Variables := {seq(seq(seq(cR[m-0-j,j,dR-q-r,q,r], q=0..dR-r), r=0..dR), j=0..m-0)}
union {seq(seq(seq(cS[m-3-j,j,dS-q-r,q,r], q=0..dS-r), r=0..dS), j=0..m-3)}
union {seq(seq(seq(cU[m-6-j,j,dU-q-r,q,r], q=0..dU-r), r=0..dU), j=0..m-6)}
union {seq(seq(seq(cV[m-9-j,j,dV-q-r,q,r], q=0..dV-r), r=0..dV), j=0..m-9)}
union {seq(seq(seq(cW[m-12-j,j,dW-q-r,q,r], q=0..dW-r), r=0..dW), j=0..m-12)}:
Nombre_Variables := nops(Variables);
Solutions := msolve(Equations,p):
Verification := seq(op(2,op(i,Solutions)),i=1..nops(Solutions)):
indets({Verification}):
nops(indets({Verification})):
assign(Solutions):
for j from 0 to m-0 do
Voir_R[m-0-j,j]:= add(add(cR[m-0-j,j,dR-q-r,q,r]*T^((dR-q-r))*X^q*Y^r, q=0..dR-r), r=0..dR):
od:
for j from 0 to m-3 do
Voir_S[m-3-j,j]:= add(add(cS[m-3-j,j,dS-q-r,q,r]*T^((dS-q-r))*X^q*Y^r, q=0..dS-r), r=0..dS):
od:
for j from 0 to m-6 do
Voir_U[m-6-j,j]:= add(add(cU[m-6-j,j,dU-q-r,q,r]*T^((dU-q-r))*X^q*Y^r, q=0..dU-r), r=0..dU):
od:
for j from 0 to m-9 do
Voir_V[m-9-j,j]:= add(add(cV[m-9-j,j,dV-q-r,q,r]*T^((dV-q-r))*X^q*Y^r, q=0..dV-r), r=0..dV):
od:
for j from 0 to m-12 do
Voir_W[m-12-j,j]:= add(add(cW[m-12-j,j,dW-q-r,q,r]*T^((dW-q-r))*X^q*Y^r, q=0..dW-r), r=0..dW):
od:
Voir_R_13,0 := 0:
Voir_R_12,1 := 0:
Voir_R_11,2 := 0:
Voir_R_10,3 := 0:
Voir_R_9,4 := 0:
Voir_R_8,5 := 0:
Voir_R_7,6 := 0:
Voir_R_6,7 := 0:
Voir_R_5,8 := 0:
Voir_R_4,9 := 0:
Voir_R_3,10 := 0:
Voir_R_2,11 := 0:
Voir_R_1,12 := 0:
Voir_R_0,13 := 0:
Voir_S_10,0 := 0:
Voir_S_9,1 := 0:
Voir_S_8,2 := 0:
Voir_S_7,3 := 0:
Voir_S_6,4 := 0:
Voir_S_5,5 := 0:
Voir_S_4,6 := 0:
Voir_S_3,7 := 0:
Voir_S_2,8 := 0:
Voir_S_1,9 := 0:
Voir_S_0,10 := 0:
Voir_U_7,0 := 0:
Voir_U_6,1 := 0:
Voir_U_5,2 := 0:
Voir_U_4,3 := 0:
Voir_U_3,4 := 0:
Voir_U_2,5 := 0:
Voir_U_1,6 := 0:
Voir_U_0,7 := 0:
Voir_V_4,0 := 0:
Voir_V_3,1 := 0:
Voir_V_2,2 := 0:
Voir_V_1,3 := 0:
Voir_V_0,4 := 0:
Voir_W_1,0 := 0:
Voir_W_0,1 := 0:
Appendix B A Cohomological Observation
This appendix was authored by Pengchao Wang
(School of Mathematical Sciences, Peking University, Beijing 100871, P. R. China).
Email: 2200010773@stu.pku.edu.cn.
Let $\mathcal{C}\subset\mathbb{P}^{2}$ be three generic smooth conics with simple normal crossings. Let
$0\subset S^{m}\,T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})=F^{0}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset F^{1}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset\cdots\subset F^{\lfloor m/3\rfloor}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))=E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$
be the ascending filtration of subsheaves that defines the grading (2.3). We set $F^{k}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))=0$ for $k\leqslant-1$. From (2.3) we obtain:
Proposition B.1.
For every $0\leqslant k\leqslant m/3$, the following sequence is exact:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{k-1}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{k}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${S^{m-3k}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\overline{K}_{\mathbb{P}^{2}}^{\otimes k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
For convenience, from now on we write $\overline{T}^{*}_{\mathbb{P}^{2}}\coloneqq T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ and $\mathcal{O}(1)\coloneqq\mathcal{O}_{\mathbb{P}^{2}}(1)$.
Proposition B.2.
For integer $q\geqslant 0$ and $r\in\{1,2\}$, the following vanishing holds:
$H^{0}\big(\mathbb{P}^{2},E_{2,3q+r}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\mathcal{O}_{\mathbb{P}^{2}}(-(3q+r))\big)=0,$
provided that $H^{0}\big(\mathbb{P}^{2},S^{3\ell+r}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\mathcal{O}_{\mathbb{P}^{2}}(-(3\ell+r))\big)=0$ for all $\ell=0,1,\dots,q$.
Proof.
By Proposition B.1, for every integer $k=0,1,\dots,q$, we have a short exact sequence:
$0\rightarrow F^{k-1}(E_{2,3q+r}\overline{T}^{*}_{\mathbb{P}^{2}})\otimes\mathcal{O}(-(3q+r))\rightarrow F^{k}(E_{2,3q+r}\overline{T}^{*}_{\mathbb{P}^{2}})\otimes\mathcal{O}(-(3q+r))\rightarrow S^{3q+r-3k}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\overline{K}_{\mathbb{P}^{2}}^{\otimes k}\otimes\mathcal{O}(-(3q+r))\rightarrow 0.$
Note that the last term $S^{3q+r-3k}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\overline{K}_{\mathbb{P}^{2}}^{\otimes k}\otimes\mathcal{O}(-(3q+r))\cong S^{3(q-k)+r}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\mathcal{O}(-(3(q-k)+r))$ has no global section by our assumption. Considering the induced long exact sequence in cohomology, we obtain
$H^{0}\big(\mathbb{P}^{2},F^{k}(E_{2,3q+r}\overline{T}^{*}_{\mathbb{P}^{2}})\otimes\mathcal{O}(-(3q+r))\big)\cong H^{0}\big(\mathbb{P}^{2},F^{k-1}(E_{2,3q+r}\overline{T}^{*}_{\mathbb{P}^{2}})\otimes\mathcal{O}(-(3q+r))\big).$
By iterating backwards for $k=q,q-1,\dots,1$, we conclude the desired vanishing since $H^{0}\big(\mathbb{P}^{2},S^{3q+r}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\mathcal{O}(-(3q+r))\big)=0$ by assumption. ∎
Remark B.3.
The above argument fails for $r=0$ because when $q=k$, the term
$S^{3q+r-3k}\overline{T}^{*}_{\mathbb{P}^{2}}\otimes\overline{K}_{\mathbb{P}^{2}}^{\otimes k}\otimes\mathcal{O}(-(3q+r))\cong\mathcal{O}_{\mathbb{P}^{2}}$
has one-dimensional global sections.